### v1

# Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries

Abstract
Retrieval-Augmented Generation (RAG) is a popular approach for enhancing Large Language Models (LLMs) by addressing their limitations in verifying facts and answering knowledge-intensive questions. As the research in LLM extends their capability to handle input modality other than text, e.g. image, several multimodal RAG benchmarks are proposed. Nonetheless, they mainly use textual knowledge bases as the primary source of evidences for augmentation. There still lack benchmarks designed to evaluate images as augmentation in RAG systems and how they leverage visual knowledge. We propose Visual-RAG, a novel Question Answering benchmark that emphasizes visual knowledge intensive questions. Unlike prior works relying on text-based evidence, Visual-RAG necessitates text-to-image retrieval and integration of relevant clue images to extract visual knowledge as evidence. With Visual-RAG, we evaluate 5 open-sourced and 3 proprietary Multimodal LLMs (MLLMs), revealing that images can serve as good evidence in RAG; however, even the SoTA models struggle with effectively extracting and utilizing visual knowledge [^1] Benchmark and evaluation codes will be available at https://github.com/LuciusLan/Visual-RAG..
Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries
Yin Wu1, Quanyu Long1, Jing Li2, Jianfei Yu3, Wenya Wang1 1 Nanyang Technological University, 2 Harbin Institute of Technology (Shenzhen), 3 Nanjing University of Science and Technology wuyi0023@e.ntu.edu.sg
1 Introduction
Figure 1: Comparison on query format and QA pipeline between InfoSeek-style and our Visual-RAG. InfoSeek-style: The entity in the query image must first be recognized to extract textual evidence, which is then used independently to answer the question. Visual-RAG: The query is text-only, but direct answers are rarely available in textual form. The model must retrieve relevant images and interpret them to generate an answer.
Benchmark Query Modality Knowledge Modality Retrieval Modality InfoSeek (Chen et al., 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] E-VQA (Mensink et al., 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] WebQA (Chang et al., 2022) [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] MRAG-Bench (Hu et al., 2024) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Visual-RAG (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Table 1: Comparing modalities of knowledge intensive VQA benchmarks. The link icon [Uncaptioned image] denotes that the two objects are paired, retrieving one will automatically link to the other. For InfoSeek and E-VQA, Yan and Xie (2024) demonstrated that by image-to-image retrieval on Wikipedia images, the Wikipedia articles containing those images are also retrieved at a significantly higher recall comparing to image-to-text retrieval. Similarly for WebQA, a question-to-caption (text-to-text) retrieval achieves better performance than text-to-image retrieval, as shown in Table 5 in Appendix. Our Visual-RAG enforces text-to-image retrieval for visual knowledge.
We have observed rapid advancement of Large Language Models (LLMs) in recent years. However, they are prone to generating hallucinated responses Ji et al. (2023). Retrieval-Augmented Generation (RAG) has emerged as a widely adopted approach to mitigate this challenge (Lewis et al., 2020; Karpukhin et al., 2020; Izacard and Grave, 2021). By retrieving and integrating relevant textual documents from external knowledge sources, it enables LLMs to answer knowledge-intensive questions more accurately. Yet, many real-world scenarios demand multimodal knowledge (e.g., visual cues or geographical maps), suggesting that text-only retrieval may be insufficient.
Building on the success of text-only RAG, various evaluation benchmarks have been proposed to assess the capability of Multimodal LLMs (MLLMs, a.k.a. Large Vision Language Models LVLM) in using multimodal RAG to address complex, multimodal knowledge intensive questions, including InfoSeek (Chen et al., 2023), Encyclopedic-VQA (Mensink et al., 2023), etc. They feature mixed-modal queries where the question is asked based on a paired image.
However, existing datasets reveal several limitations. Firstly, although images are included in the query, MLLMs do not utilize the visual knowledge from these images when generating answers. Secondly, questions in these datasets typically do not require reasoning over knowledge extracted from images. Figure 1 gives an example on InfoSeek-styled query. To answer the query, a model needs to first recognize the entity in query image, then locate the Wikipedia article and augment relevant passage to LLMs to generate an answer. Query image information is only required in the entity recognition step. These existing multimodal benchmarks are textual knowledge intensive, highlighting the need for a new benchmark for visual knowledge intensive QA.
To this end, We present Visual-RAG, a visual knowledge intensive QA benchmark designed to evaluate text-to-image retrieval and answer generation augmented by retrieved images. Key features of Visual-RAG are:
  1. 1.
    The entities in queries are obscure, with their descriptions being brief, rarely documented in standard references (e.g., Wikipedia)
  2. 2.
    The visual features queried are fine-grained, with only 5.35% of the images in the corpus containing the specific visual evidence needed to answer the questions, posing a challenging text-to-image retrieval task.
  3. 3.
    The majority of non-clue images are hard-negatives which depict the same entities and are visually similar. This subtle distinction further increases the difficulty of cross-modal retrieval.
To construct Visual-RAG, we begin with LLM-generated candidate queries, which are then refined through human filtering and rewriting. Next, we employ open-sourced MLLM and human annotators to perform another filtering step to ensure an appropriate level of difficulty for the queries. Detailed procedures for the entire benchmark construction can be found in Section 4.
With Visual-RAG, we comprehensively evaluate 8 mainstream MLLMs, including 5 open-sourced models and 3 proprietary ones. Firstly, we demonstrate that images can serve as powerful evidence for augmented generation through baseline experiments; secondly, we experiment with various RAG settings, testing the ability of MLLMs digesting images as augmentation, and analyse model behaviour on identifying clue from irrelevant images. Our benchmark sheds light on real-world scenarios for evaluating MLLMs, when users raise natural language queries for knowledge-intensive questions where textual evidence is scarce but visual evidence possibly exists, posing the challenging task of retrieving and utilizing relevant images in the wild.
The key findings from our experiments are summarized as follows:
  • Cross-modal retrieval is challenging. The commonly used small cross-modal retriever (CLIP) faces significant challenges addressing difficult queries requiring the identification of fine-grained visual features.
  • The amount of evidence affects retrieval-augmented generation. The evaluated MLLMs demonstrate the ability to extract visual knowledge to answer questions. However, a single ground-truth clue image is insufficient for the models to achieve optimal performance.
  • Negatives can be helpful. Providing multiple images as input, including both clue images and non-clue images, could enhance performance for proprietary models. This improvement likely stems from the models’ ability to contrast clue and non-clue images, enabling them to better process the relationship between the question and the images.
  • Inverted dynamics of open-sourced and proprietary models. Open-source models excel with a single clue image but fail to differentiate the clue when irrelevant images are also introduced; while proprietary models are initially conservative with a single clue but ultimately achieve better performance through multi-image augmentation. This reveals open-source models are weaker at handling multiple images, compared to proprietary models.
2 Related Work
While numerous benchmarks evaluate the capabilities of MLLMs, we focus here on those designed for knowledge-intensive visual question answering (VQA), where RAG is the primary solution paradigm. Table 1 compares the modalities covered by commonly used benchmarks. A more detailed discussion of how Visual-RAG differs from prior work is provided in Appendix A.
Textual Knowledge Centric VQA Benchmarks
OK-VQA (Marino et al., 2019) and its augmented version A-OKVQA (Schwenk et al., 2022) focus on commonsense or world knowledge, though they do not strongly depend on external knowledge retrieval. ViQuAE (Lerner et al., 2022), InfoSeek (Chen et al., 2023) and Encyclopedic-VQA (E-VQA, Mensink et al., 2023) address “encyclopedic” questions, paired with query images of specific entities, utilizing information retrieved from external textual sources (e.g., Wikipedia) to answer.
Visual Knowledge Centric QA Benchmarks
While the above benchmarks all position themselves as knowledge-intensive VQA tasks, it is important to note that the required knowledge remains primarily textual, with images functioning mostly as entity anchors. In contrast, to our knowledge, only two related works truly require image retrieval as evidence for answering questions.
WebQA (Chang et al., 2022) is an earlier effort that requires retrieving images for visual knowledge. However, it heavily relies on captions because the images themselves usually lack indicators of which are clue images. A concurrent work, MRAG-Bench (Hu et al., 2024) shares a similar motivation in using image retrieval for question answering. However, it primarily aims to recognize an entity when the instance in query image is incomplete or partially obscured, relying on retrieved images that present a complete view of the same entity. This entity-recognition objective differs from our focus on answering visual-feature-oriented factual queries. As queries in MRAG-Bench include image, the model can use the “shortcut” of matching a visually similar retrieved candidate, bypassing the need for deeper text-to-image reasoning.
3 Benchmark Overview
We propose Visual-RAG, a text-only Question Answering benchmark for challenging, visual knowledge intensive factual questions. It is an evaluation benchmark for visual knowledge centric text-to-image retrieval, as well as a benchmark for evaluating MLLMs’ abilities on extracting visual knowledge as augmentation under retrieval-augmented generation settings.
Visual Knowledge Intensive Queries
Our Visual-RAG focuses on text-only queries that probe factual knowledge of visual feature which generally hold true of an entity (e.g. “What colour are the stamens of Chamaecrista nictitans”), rather than instance-specific attributes commonly addressed in standard VQA tasks (e.g. “What colour is the cup on table”). Hence, we specifically select the organism domain which enables specialized queries. Unlike existing works, we deliberately exclude images from the queries due to the reason that including an image often reduces the problem to mere entity recognition or unimodal-similarity match, instead of our main objective of cross-modal knowledge extraction. Numerous established benchmarks already cover entity recognition, including the OVEN (Hu et al., 2023) dataset which serves as the foundation for the the InfoSeek dataset, and the iNaturalist 2021 (Van Horn et al., 2021) dataset, which we employ in this work. Moreover, text-only queries better align with real-world scenarios where users tend to raise questions about an entity without an image.
Naturally Co-occurring Hard Negative Images
For each question, multiple clue images might be available, reflecting real-world scenarios. Additionally, the dataset incorporates a substantial number of naturally co-occurring hard negative images, i.e., images of the same species that do not exhibit the visual feature in query. Notably, all clue images for a given question can be retrieved purely based on visual information, without reliance on captions.
Dataset Statistics
The current version of Visual-RAG contains 400 queries, forming an image knowledge base of total 103824 images. The average clue image count for each query is 13.89, and the clue images occupy 5.35% of the image knowledge base. The distribution of organism categories and query types is shown in Figure 2.
(a) Distribution of organism categories in Visual-RAG
(b) Distribution of question categories
Figure 2: Distribution of organisms and question categories.
4 Benchmark Construction
4.1 Data source
Following Encyclopedic VQA Mensink et al. (2023), we adopt iNaturalist 2021 (iNat21, Van Horn et al., 2021) as the cornerstone of our benchmark. iNat21 contains 2.7M images of organism (animals, plants, fungi, etc.), spanning across 10000 taxonomy species. Each species in iNat21 has 200–300 images, originally collected for species-level image classification.
For organism species selection in query construction, we aim to ensure that the chosen species are relatively underrepresented in textual resources. To achieve that, we filter them based on the length of their Wikipedia summary and description sections. Well-known species typically have multiple sections (e.g., habitats, behaviour, etc.), indicating extensive documentation. We therefore include only those species whose summary and description account for no more than 50% length of the entire Wikipedia article. This also helps minimize reliance on prior knowledge: widely recognized species may have detailed descriptions on visual features across LLM pre-training corpora, even if these details are not explicitly mentioned in Wikipedia.
4.2 Query Collection
Question generation.
Our objective is to create queries about distinct visual features of organisms, ensuring that only a small fraction of images within each species display the queried feature. Each question must yield a unique answer and generally hold true for every instance of that species, preserving both specificity and accuracy. Typically, composing such specialized queries would require domain expertise and a labour-intensive review of the 200–300 images per species.
Consequently, we employ OpenAI-o1 (OpenAI, 2024) for query generation to replace human experts. As creating meaningful questions requires expert knowledge, we provide the Wikipedia summary and description sections for the model to understand the species. We explicitly instruct the model NOT to generate questions about properties already covered in these text passages, prompting it to focus on more specialized visual attributes unlikely to appear in standard references. The reasoning and self-validating ability of o1 enables it to follow the complex instructions, leading to high-quality and diverse prototype queries. The model is tasked with generating both a set of candidate queries and their corresponding visual features. A refined human annotator filtering and rewriting process is then applied based on these visual features, to ensure query validity [^2] More details about the query generation prompt and human filtering process can be found in Appendix C, D..
Question Filtering and Image Annotation.
The question generation process cannot guarantee compliance with the previously mentioned requirement, i.e., clue images containing the answer should remain a minority in the set; also, query should not be so difficult that no clue image exists in the corpus. We designate a threshold clue image rate, that at most 20% of images in a species level sub-corpus can be clue.
Firstly, we employ an open-sourced MLLM [^3] Llama-3.2-11B-Vision-Instruct was used. Prompt template can be found in Appendix D as a coarse-level filter. For each query, the model iterates over the sub-corpus one image at a time, binary labelling the images by whether the referenced feature is present. These labels are used to roughly filter out queries that exceed the 20% clue-image threshold. Our rationale is that if an unusually large fraction of images exhibit a visual feature, it is likely prominent for the MLLM to detect and generate labels that are reliable enough.
Secondly, we recruit volunteer university students as annotators to label the images. These annotators identify whether each image provides evidence for the question and thereby confirm whether (a) at least one valid clue image exists, excluding overly difficult queries having no clue image and (b) fewer than 20% of the images are clue, excluding remaining “easy” queries not recognized by the coarse-filter. At the same time, the annotators provide the ground-truth answers for each query. This two-stage process ensures our final dataset meets both the clue-image minority requirement and the fundamental criterion that each query should be indeed answerable by evidence clue image within the corpus.
In the current version of the benchmark, we initially select around 1000 questions that passed the coarse filter. After human annotation, 400 of these questions satisfy the clue-image rate threshold and are confirmed answerable with at least one valid clue image.
5 Experimental Setup
5.1 Retrieval Evaluation
For retrieval, we evaluate the performance with Normalized Discounted Cumulative Gain (NDCG) and hit rate – the probability that at least one clue image is found among the top-$k$ retrieved samples (NDCG@k and hit@k).
Due to the challenge of complex domain-specific queries, it is significantly difficult for the retrievers to search through the entire iNat21 corpus (2.7M images) while the query is about a sophisticated feature for a single species. To enhance retrieval effectiveness and emphasize the more fine-grained and distinguishing features of a species, we narrow the search space by limiting the retrieval corpus to a set of 200–300 images corresponding to the species for each query. In all following experiments, we adopt the commonly used retriever clip-vit-large-patch14-336 (Radford et al., 2021) for image retrieval. The search indexes are constructed with the FAISS library (Douze et al., 2024) using flat inner product index.
5.2 Generation Evaluation
Previous work has relied predominantly on exact-match or recall-based metrics to assess answer correctness. However, such methods are prone to overlooking partial hallucinations blended with ground-truth answers. For example, if the true answer is “black and white” but the model predicts “black and white with yellow dots,” the spurious mention of “yellow dots” remains undetected under exact-match criteria. To address this limitation, we employ GPT-4o as a more nuanced evaluator. Detailed evaluation prompt template can be found in Appendix D. A series of works, including but not limited to Kamalloo et al. (2023); Zheng et al. (2023); Huang et al. (2024), have demonstrated that general-purpose LLMs can reliably judge open-ended QA responses. We also manually validate the automatic evaluation results on a sample of 100 predicted answers, achieving a 94% accuracy rate.
In addition, we report the ROUGE score (Lin, 2004), a metric commonly used in summarization evaluation, as a relaxed version of exact match score. The ROUGE scores are omitted in result analysis section, and can be found in Appendix E.
5.3 Multimodal LLMs
We evaluate 3 advanced proprietary models: GPT-4o (OpenAI, 2025), Gemini-1.5-Pro (Google, 2025), Claude-3.5-Sonnet (Anthropic, 2025); and 5 recent open-sourced models that can process multiple images, of sizes ranging from 4B to 12B: Phi-3.5-vision-instruct (4B, Microsoft, 2024), Qwen2-VL-7B-Instruct (Wang et al., 2024), InternVL-2.5-8B (Chen et al., 2024), Llama-3.2-11B-Vision-Instruct (Meta Llama Team, 2024), and Pixtral-12B (Mistral AI, 2024). We use the default setting given by each model to perform question answering.
6 Results
6.1 Text-to-Image Retrieval
Metric @1 @5 @10 @20 @30
NDCG 26.06 26.46 30.67 38.50 45.80
Hit Rate 26.06 52.73 64.55 72.73 77.87
Hit Count 0.26 1.12 2.08 3.60 4.85
Table 2: Retrieval Results using the sub-corpus of 200-300 images for each query. Even within the small species level corpus, the model struggles with our challenging text-to-image retrieval task.
Successfully retrieving clue images that contain the required visual knowledge is critical to the overall performance of RAG systems in answering the questions. The hit rate highlights the difficulty of text-to-image retrieval in our Visual-RAG benchmark. Even within the small sub-corpus of 200 - 300 images, there is still 22% chance that no clue image can be retrieved within the top-30 samples. The NDCG score also indicates the challenge of this task, i.e., the retrieved clue images are not assigned top similarity scores comparing to the hard negative images.
6.2 Retrieval-Augmented Generation
Model Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
Baselines
Zero-shot 27.57 32.04 26.17 30.21 35.80 47.21 42.78 42.43
Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.19
Non-clue 27.77 33.27 31.53 34.94 31.63 16.61 6.25 8.92
Retrieved Top-K
K=1 33.47 39.24 34.13 37.45 34.59 28.09 19.42 18.53
3 31.14 38.05 32.93 39.44 36.72 35.26 45.48 28.49
5 35.66 40.24 35.52 37.25 39.51 40.84 45.35 30.88
7 33.67 40.44 35.33 37.25 36.65 43.82 46.81 35.86
10 33.47 39.04 37.18 36.85 37.18 44.02 48.40 38.84
15 33.86 39.64 34.39 37.25 35.99 44.42 51.20 45.62
20 32.27 40.84 36.59 - - 47.61 53.19 42.23
Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.79
One-in-K
K=3 33.03 40.72 36.72 37.21 37.00 43.90 45.64 32.87
5 33.97 39.68 35.22 39.70 35.92 45.90 46.35 35.98
7 33.35 39.36 34.80 35.66 36.53 45.31 45.46 36.60
10 32.27 39.16 35.61 36.57 34.51 46.14 46.90 37.63
15 32.23 39.40 34.99 37.32 34.59 46.54 46.58 37.86
20 32.39 37.73 34.08 - - 45.62 49.57 35.43
Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.79
Table 3: Main experiment results. The coloured cells shows the difference with oracle score, best performing $k$ values for each model is highlighted with boldface.
Does image as augmentation benefit MLLMs?
To assess whether image can be leveraged as evidence in the RAG system, we conduct baseline experiments and compare performance under three conditions: 1) Zero-shot: directly prompting MLLMs with questions without retrieval; 2) Oracle: augmenting the question with a gold clue image; 3) Non-clue: augmenting each question with a non-clue image within the same species, which does not contain visual knowledge relevant to the question [^4] For 2) and 3), the results reported are the average of 5 runs with different clue / non-clue images selected as augmentation., as shown in the first section of Table 3.
All five open-sourced models exhibit performance gains of up to 15 points when using oracle image as an augmentation. This indicates the necessity of including relevant visual knowledge to answer the questions. We conclude that for open-sourced models, a single relevant image can serve as good evidence to enhance RAG performance.
Whereas interestingly, oracle performance for proprietary models drops below the zero-shot setting. We hypothesize that the single-image clue provides insufficient evidence to establish the correct visual relationship, causing the model to respond conservatively with “I cannot decide”. These conservative responses count as failing to produce an answer, thus reducing the overall scores. Specifically, such conservative response occurs in 15.8% of samples for GPT-4o, 7.4% for Gemini, and 16.4% for Claude. Meanwhile, results for the Non-clue setting indicate that the models effectively identify these non-clue images as not containing visual knowledge. Specifically, all proprietary models frequently respond with “I don’t know” with non-clue images, leading to a substantial drop in scores comparing to the zero-shot and oracle setting.
Can MLLMs distinguish the clue from retrieved hard negatives?
We have shown that a relevant clue image can serve as good evidence in augmentation. However, retrieval performance for these challenging queries is far from perfect, as shown in the previous section. For RAG system in real implementation where a number of possibly irrelevant images are retrieved, does retrieval still help the final QA performance?
To explore this realistic RAG scenario, we provide the top-$k$ retrieved images, derived from Section 6.1, as augmentation to the MLLMs [^5] For pixtral and Llama3.2-V, $k=20$ was not evaluated due to hardware limitation.. The retrieved images are sorted in descending order based on their similarity scores. As shown in the second section of Table 3, the open-source models generally lag behind their single-oracle scores, indicating limited multi-image processing capabilities. Most open-source models struggle to isolate the clue image and effectively leverage its visual information when it is mingled with hard negatives.
In contrast, all three proprietary models exhibit a generally monotonic increase in performance as $k$ increases, except Claude which decreases at $k=20$. Notably, the performance at $k=5$ begins to surpass oracle baselines, continuing to improve along with higher $k$ values. We attribute this improvement to the comparative evaluation enabled by providing both clue and non-clue images. Compared to a single clue image, the top-$k$ RAG setting allows the models to contrast relevant and irrelevant visual features, leading to greater confidence in extracting knowledge from the clue images, whereas a single clue image may be insufficient to fully establish its relationship to the question. However, the proprietary models fail to exceed their zero-shot scores until fairly large $k$ values, suggesting that interpreting complex visual cues remains challenging, and they may even struggle with reconciling wrongly interpreted visual knowledge with their prior knowledge gained through pre-training.
As indicated by the retrieval results, even at $k=20$, there is 27% chance that no clue image is included in the retrieved set. This likely explains the low QA performance at smaller $k$ values, where failing to retrieve the clue image entirely negates any benefit of visual augmentation. To investigate further, we conduct an additional experiment, constructing image sets in which exactly one clue image and $k-1$ non-clue images are provided [^6] Both clue and non-clue images were randomly selected across 5 different combinations, results reported are the average of 5 runs for each $k$.. The clue image is positioned at the beginning of the $k$ images to simulate the perfect retrieval scenario.
Demonstrated in the last section of Table 3, the open-sourced models now exhibit decreasing performance as $k$ increases, further strengthen the conclusion of their limited capacity for handling multiple images simultaneously. By contrast, all three proprietary models surpass their single-image oracle baseline from $k=3$ onward and maintain near-constant performance at higher $k$ values. This suggests that proprietary models are capable of isolating the clue image from irrelevant images, regardless of the number of input images [^7] Additional experiments on randomly positioning the clue image within the set is also carried out. The proprietary models keeps the flat performance curve against $k$, indicating there is no positional bias..
Model Zeroshot Score Augmenting Text
Phi3.5-V 27.57 26.02
Qwen2VL 32.04 25.61
InternVL2.5 26.17 19.70
Pixtral 30.21 19.74
Llama3.2-V 35.80 29.73
Table 4: Results when augmenting retrieved Wikipedia text as evidence.
6.3 Ablation
Can the answers for Visual-RAG be found in Wikipedia?
We construct a small textual knowledge base from Wikipedia articles for the 10k species in iNat21 dataset. This minimum knowledge base comprising approximately 97,000 chunks, each consisting of $\sim$150 tokens. The detailed collection process is described in Appendix B. For text retrieval, we use a pre-trained dense text retriever, DRAGON Plus (Lin et al., 2023). Given that the ground-truth labels for relevant text chunks are not available, we only assess whether the retrieved text chunks correspond to the correct Wikipedia page, and the retriever achieves HitRate@10 of 97%. Next, we present the QA performance across open-sourced models augmented with top-10 retrieved textual knowledge chunks. As shown in Table 4, augmenting the question with textual knowledge from Wikipedia does not improve answer quality for most models. Instead, in numerous cases, models fail to locate relevant information within the augmented text and output “I don’t know”. This behaviour in turn leads to scores lower than the zero-shot setting.
These experiments provide strong evidence that the knowledge required to answer the questions related to visual features in our benchmark is unlikely to be contained in textual knowledge bases.
Is More Clues Always Better?
We observed that for proprietary models, top-$k$ scores at higher $k$ values excel their 1-in-$k$ score. Similarly, the open-source models also achieve their best performance at larger $k$ value in top-$k$ setting than in 1-in-$k$. A likely explanation is that the expected number of clue images retrieved in top-$k$ is substantially higher than when only a single clue image is guaranteed. We in turn ablate with effect of multiple clue images. Fixing $k=10$, we experiment with varying numbers of clue images with $n\in\{1,2,3,4,5\}$ ($n$-in-10). We select queries that contain at least 5 clues to enable this experiment. As shown in Figure 3, all models exhibit increasing performance as $n$ grows, aligning with the intuition that additional clue images provide more robust visual evidence.
Figure 3: Impact from number of clue images
Hard-negatives can be helpful; how about easy-negatives?
As an ablation to the previous experiment on whether multiple images benefit the overall QA performance, we now mix one clue image with $k-1$ highly irrelevant images from subsets that are not the query species. The results indicate that both open-source and proprietary models benefit more from hard negatives coming from the same species compared with easy negatives coming from different species. This further strengthen our conclusion that hard negatives are helpful in contrasting the queried features to inform the correct answer. Detailed results can be found in Table 9 in Appendix.
7 Conclusion
In this paper, we present Visual-RAG, a challenging VQA benchmark that requires visual knowledge as evidence to answer questions. Different from previous works that include images as part of queries and let models retrieve textual knowledge, our Visual-RAG is formed with text-only queries and enforces text-to-image retrieval for challenging visual knowledge. We evaluate 8 popular MLLMs under various multimodal RAG settings. The results indicate that images can serve as knowledge sources for answering visual knowledge intensive questions, but the models struggle to effectively extract visual knowledge from a single clue image. Surprisingly, the current models are able to reason across clue and non-clue images for better understanding of the questions. As one of the pioneer works in visual-knowledge centric multimodal RAG, our benchmark can help promote the research and development of MLLMs in extracting fine-grained visual knowledge and enhance cross-modal retrieval in a more realistic setting.
Limitations
While our Visual-RAG addresses the lack of visual knowledge intensive QA benchmarks, it has several design constraints. First, our query generation and manual rewriting process focus on single-hop questions only. Although a small subset of queries might involve simple comparisons within the same image, such as checking the colour of a bird’s upper beak vs. lower beak, there are no truly multi-hop queries requiring reasoning across multiple images. Designing such multi-image or even cross-species queries would demand significantly more expert knowledge, and we leave this as an avenue for future work.
Secondly, Visual-RAG focuses exclusively on organisms, which raises questions about the broader generalizability of our findings. We chose the domain of organisms because they allow for factual queries regarding visual features that remain consistent across multiple individuals of a species, yet are present in only a minority of images. Other commonly used domains in knowledge-intensive VQA such as buildings, artworks, news event, etc., do not always meet these criteria, as answers may not hold universally for all instances or there exists only one instance, and the required visual features are often too prevalent or too trivial to foster a genuinely challenging retrieval scenario.
Lastly, the overall scale of Visual-RAG is limited compared to larger benchmarks like InfoSeek and Encyclopedic-VQA. Their heuristic-driven methods for automatically generating queries and answers do not readily apply to our specialized setting, which hinges on expert-informed checks to ensure the presence (or absence) of specific visual features. Despite these constraints, Visual-RAG represents an important step toward truly visual knowledge intensive QA, and we hope it will inspire further research into more complex, multi-hop, and cross-domain tasks.
Ethical Considerations
Licensing
We do not own, nor do we redistribute the images of iNaturalist 2021 dataset. Visual-RAG only provides the annotations of queries and answers, as well as image labels on whether it is clue to query, linked to image GUID in iNatrualist 2021 dataset. All images in iNat21 dataset are shared under one of the following Creative Commons licenses, in particular: CC BY 4.0, CC BY-NC 4.0, CC BY-NC-ND 4.0, CC BY-NC-SA 4.0, CC0 1.0, CC BY-ND 4.0, CC BY-SA 4.0. Any usage of the images is subject to iNaturalist and iNaturalist 2021 dataset’s term of use. We strictly adhere to the intended non-commercial research use granted by the iNaturalist community and iNaturalist 2021 dataset, and we share our annotation in Visual-RAG under CC BY-NC 4.0 license. The authors assume no responsibility for any non-compliant usage or legal/ethical issues related to the original iNaturalist 2021 dataset.
Potential Risks
The annotation in this dataset, i.e. the queries, answers and image labels, are intended for research purposes only. Our dataset’s queries, answers, and image labels are provided by volunteer annotators who are not professional taxonomists or biologists. We make every effort to ensure correctness, such as proactively consulting publicly available expertised resources of biology and taxonomy over internet, and annotators are undergraduate students who possess sufficient English proficiency to utilize such publicly available resources for fact validation. Though, we cannot guarantee the absence of factual errors. We make no guarantees regarding completeness or correctness for real-world decision-making.
Our annotation does not contain any personally identifiable information, nor does it reference private data, locations, or individuals. Our annotation does not reveal exact geographical coordinates or sensitive ecological information that might pose risks to endangered species or protected habitats. For other potential risks regarding images in iNaturalist 2021 dataset, please refer to the original paper and iNaturalist website (www.inaturalist.org).
Appendix A Relation with previous works
Knowledge intensive VQA benchmarks focusing textual knowledge, including but not limited to: ViQuAE (Lerner et al., 2022), InfoSeek (Chen et al., 2023) and Encyclopedic-VQA (Mensink et al., 2023), are discussed in related works. These benchmarks contain questions paired with images of certain entities, and knowledge for such entities can be found in an external textual knowledge base (e.g., Wikipedia). The questions are deliberately constructed to omit the entity names (e.g., “How many feet tall does the plant grow to?” instead of “How many feet tall does Acacia paradoxa grow to?”).
Typically, given a question and its associated image, a model is expected to: 1) identify the entity depicted in the image, 2) retrieve relevant documents or passages from a knowledge base that pertain to both the question and the entity, and 3) generate an answer based on the question, image, and retrieved documents. Step 1) and 2) is usually carried out by leveraging vision-language pretrained encoders, such as CLIP (Radford et al., 2021), to facilitate cross-modal (image-to-text) retrieval.
While the above mentioned benchmarks are all claimed to be Knowledge-Intensive VQA benchmarks, it is noteworthy that the knowledge to be retrieved is textual knowledge. The image is only served as an anchor of the entity. While cross-modal retrieval was commonly used, the knowledge retrieval process can actually be achieved in uni-modal manner. A recent work (Yan and Xie, 2024) achieved “SoTA” performance on InfoSeek and E-VQA employing image-to-image retrieval within a knowledge corpus comprising Wikipedia pages and their associated images. Entity recognition is performed by identifying the most visually similar Wikipedia image to the query image, which automatically enables retrieval of the related knowledge document (i.e., the Wikipedia page containing that image). Answers are then generated using a text-only LLM, augmented with the relevant paragraph from the Wikipedia page, which is extracted through fine-grained reranking.
WebQA dataset (Chang et al., 2022) included questions that require retrieving images and utilizing visual knowledge to generate answers. However, its retrieval process is heavily reliant on image captions, as the images themselves typically lack indicators identifying them as clue images for answering a question. For instance, considering the question: “Are the land dinosaurs guarded by rail in both the Display Museum of Natural History in University of Michigan and the Museo Jurassic de Asturias?” Without captions, the model would be unable to discern which dinosaur fossil image corresponds to which museum.
Table 5 shows the retrieval scores in our implementation of retrieval experiments on WebQA benchmark for queries having images as knowledge source. While the queries are text-only, aligning with our setting, the text-to-image retrieval scores significantly lag behind text-to-caption retrieval, especially for lower $k$.
Recall @1 @5 @10 @20 @50
Text-to-Image 15.17 37.08 47.63 58.43 72.32
Text-to-Caption 43.25 68.74 76.86 83.79 89.92
Table 5: Retrieval scores for WebQA dataset.
MRAG-Bench (Hu et al., 2024) shares a similar objective to ours by proposing a benchmark that requires retrieving visual knowledge to solve questions. However, despite this similarity, the focus remains on entity recognition. Most tasks are structured such that an image displaying an incomplete or partially visible entity is provided, and the retrieval system is tasked with identifying images that depicts the entire entity. This approach primarily enhances the MLLM’s ability to recognize entities rather than addressing broader knowledge-intensive tasks. Furthermore, non-clue images in MRAG-Bench are from different entities, which are visually distinct, in contrast to the numerous hard negatives in our Visual-RAG .
Concurrent to our work, the iNaturalist team released 2024 version of their natural image dataset, iNat24 (Vendrow et al., 2024), coming with a text-to-image retrieval benchmark named INQUIRE, containing 250 queries. Though, the queries in INQUIRE are designated for image retrieval only, and are not QA style questions that have explicit answers. As iNat24 and iNat21 share the same taxonomy classes of organisms, we do consider adding the new images in iNat24 to our benchmark.
We also note a related line of research (Tanaka et al., 2023; Van Landeghem et al., 2023) on extracting knowledge from document images (e.g., PDF files, Powerpoint slides), conventionally referred to as OCR-based VQA. Although these tasks also require models to retrieve relevant page and interpret images to answer textual queries, the underlying knowledge is primarily textual—obtained via OCR. This focus diverges considerably from our goal of visual feature extraction, where visual cues are inherently sparser in the wild.
Appendix B Construction of Minimum Textual Knowledge Base
Among the 10000 taxonomy species in iNat21 dataset, we matched 9614 Wikipedia pages. iNat21 provides the scientific name as well as one common name of the species. We conduct multiple rounds of matching: initially, Wikipedia articles are matched based on species scientific names with Wikipedia titles; for unmatched species, match their common names with article titles; further unmatched species will go through matching within article content. All articles containing the scientific name are collected, sorting according to the position of the scientific name’s first occurrence within the article. The article with earliest occurrence is selected as matched page. Due to the fact that taxonomy study frequently change the attribution of a species, the scientific name in iNat21 dataset can be deprecated and no matched page can be found for them.
We then applied basic text chunking to these articles, splitting them into chunks of approximately 150 tokens. Each chunk is bounded within a single section of the original article to preserve contextual integrity. Sections such as “Reference”, “See also”, “External links”, etc., which lack informative content, were excluded. The chunks are cut upon period mark, so that they won’t end or start with incomplete sentence. Two adjacent sections, having total length less than 150 tokens will be merged, with the section titles kept.
Appendix C Annotation details
For the query generation, we mentioned that we have gone through a human annotator originated refined filtering and rewriting process. The prompt template is shown in Table LABEL:tab:prompt_qgen, and following shows the example prototype queries generated with the sample prompt:
  1. 1.
    Does rescuegrass (scientific name: Bromus catharticus) have a visible ligule when the leaf sheath is gently parted, and if so, what is its shape? (Explanation: The ligule is a thin membrane at the junction of leaf blade and sheath, typically hidden in standard photos) (Feature: shape of ligule) Dropped. Too detailed and hard to distinguish
  2. 2.
    During the flowering stage of rescuegrass (scientific name: Bromus catharticus), what color do the exposed anthers display? (Explanation: The anthers are only briefly visible and require a closer or angled view) (Feature: color of anthers) Passed. Annotator found a image of flower anther (Figure 4) and confirm it is rare image.
  3. 3.
    Is there any distinct pattern or texture on the backside of the spikelets of rescuegrass (scientific name: Bromus catharticus), observable only when the spikelets are lifted or spread apart? (Feature: pattern on backside of spikelets) Dropped. Viewing the spikelets image, there is no “backside” on spikelets as they are symmetric.
  4. 4.
    When the plant is fully mature, how does the collar region (the area where the leaf blade meets the sheath) of rescuegrass (scientific name: Bromus catharticus) appear from the underside view? (Explanation: This area may be covered by overlapping leaves in typical photos) (Feature: shape/appearance of collar region) Dropped. Query too obscure and likely no clue image can be found.
Figure 4: Image of Bromus catharticus flower anther. Owned by original uploader at www.inaturalist.org, shared under CC-BY-NC license.
Human annotator carefully check the candidate prototype queries by browsing all available material online with the scientific name as search keyword, as well as briefly scan through the image corpus by viewing the image thumbnails.
After query generation and rewriting, and coarse-level filtering described in section 4, we dispatch the validated queries to the next round of human annotation of image labels. Figure 5 shows the annotation interface for image labelling. [^8] Similar interface is used for query filtering and rewriting annotator to browse, except the image matrix is more dense and displaying only thumbnail sized images. Annotators are required to only label “Y”, to indicate that this image contain the queried feature and is a clue to the query. Leaving the checkbox blank will be considered as negative label. This design aims to save annotators’ effort, as clue images are assumed to be the minority. The rough clue image rate predicted by MLLM in the previous step is also provided (“Y_Rate: 0.003” after the question text) as reference for annotators to estimate the query difficulty. In the meanwhile, annotators provide query answers upon finding a clue image, as shown in Figure 6. Annotators are allowed to skip for a query when finding more than 9 clues after the first 27 images, indicating this query likly fail the clue image threshold. The order of images are randomly shuffled.
Figure 5: Annotation interface for image labelling
Figure 6: Annotation interface for image label and query answer
Appendix D Prompt Template
This section list the prompt templates used through out our experiments.
Prompt Template for Query Generation
Given an organism’s scientific taxonomy name and common name (if available), please generate some questions regarding its visual feature, like color, shape, pattern, texture, etc. Please limit the question asked to be regarding features that may not be usually found in photos. This can be regarding the following but not limited to: uncommon-to-see parts of animals or plants / parts that can only be viewed at specific posture or require photo taken at specific angle / less commonly seen stage of life cycle of incect, etc. (e.g. “What is the color of paw pad of giant panda”, photos showing the paw pad of giant panda would be only a small portion; “What pattern is on the back of brown peacock’s larva?”, asking for the larva form of this butterfly.) Questions regarding non-visual feature should not be asked. (e.g., “how long can giant panda live”, “what does puma eat” are not suitable questions, as such information cannot be extracted from a photo of that animal). Please avoid questions regarding exact length/area (of tails, ears, leaf, etc.), as such attribute is hard to be estimated from photo. Please avoid questions regarding eggs, nests, nets, etc., and focus on the organism’s body. Please avoid generating multiple paraphrased questions regarding a same feature or body part. Please avoid questions comparing another species / form, but you can compare within the same individual (e.g. Does the upper beak or lower beak of XXX bird have darker color?).
Except these constraints, try to generate “interesting” questions: for example, photos showing underside of leaf may be rare, but asking what color is underside of leaf is not interesting – it’s green anyway. Hence after generating each question, try to check if the questions is “interesting” or not. Please avoid asking for features that “technically” appear on a photo but hard to see (e.g., “What color is the tiny hair on the surface of XYZ plant’s stem?” Any photo displays the stem will show the tiny hair, but it needs to be zoomed in / photo taken at very close, to see. The desired questions is regarding features that is hidden in some photos, not features that appear in every photo but simply too small to see).
A likely related paragraph from Wikipedia is provided for you to better understand the organism. Please avoid question regarding visual features explicitly mentioned in the Wikipedia paragraphs. (e.g., paragraph: “The seeds are colored in red” -> Don’t ask “What color is the seed”; paragraph: “The plant have symbolic seed pods” (but doesn’t mention its color) -> You may ask “What shape/color is the seed pod”)
Following are format instructions. You don’t need to give answer to the questions.Please generate questions by asking the common name and add scientific name in brackets, and indicate the questioned visual feature after the question sentence, e.g.: “What color is *** of {common_name} (scientific name: {scientific_name})? (Feature: color of ***)”; if there is no common name, simply ask with the scientific name. If explanation to the question is needed (such as explain a technical term), do not start a new line, append the explanation after the question enclosed by small brackets ().
Following are a few examples:
Scientific name: Diastictis fracturalis
Common name: fractured western snout moth
Question:
1. What pattern, if any, can be observed on fractured western snout moth’s (scientific name: Diastictis fracturalis) underside of wing during flying? (Explanation: the moth’s wing usually cover its whole abdomen when resting, and underside will be less commonly appear in photos) (Feature: pattern on underside of wing during flying)
2. What is the common color of larva fractured western snout moth (scientific name: Diastictis fracturalis)? (Feature: color of larva)?
Scientific name: Tortula muralis
Common name: wall-screw moss
Question:
1. What is the common color of tip of leaf of wall-screw moss (scientific name: Tortula muralis)? (Explanation: The moss can be distinguished from similar mosses by its ḧair-pointedl̈eaves.) (Feature: color of leaf tip)
Scientific name: Ptiliogonys cinereus
Common name: Gray Silky-flycatcher
Question:
1. What color are the undertail covert (short plumages that cover the long tail feather) of Gray Silky-flycatcher (scientific name: Ptiliogonys cinereus) when viewing from front? (Feature: color of undertail covert viewing from front)
2. What pattern can be observed on the tail feather of Gray Silky-flycatcher (scientific name: Ptiliogonys cinereus) from dorsal (back) view? (Feature: pattern on tail feather from dorsal view)
More ICL Examples omitted
End of examples.
Now, please generate questions for the given scientific name.
Example new query
Scientific name: Bromus catharticus
Common Name: Rescuegrass
Wiki Paragraph:
Bromus catharticus is a species of brome grass known by the common names rescuegrass, grazing brome, prairie grass, and Schrader’s bromegrass. The specific epithet catharticus is Latin, meaning cathartic. The common name rescuegrass refers to the ability of the grass to provide forage after harsh droughts or severe winters. The grass has a diploid number of 42 …… Omitted due to space. Actual prompt includes whole description passage.
Question:
Table 6: Prompt Template for Query Generation. Texts in italic and purple are not part of prompt.
Prompt Template for Main Question Answering Experiments of Different Settings
Zero-shot
Please answer the question regarding a visual feature of an organism (animal, plant, etc.). Please follow the answer format: “Answer: {answer_text}”
Question:
Oracle (one clue)
Please answer the question regarding a visual feature of an organism (animal, plant, etc.). You will be provided with a image regarding that organism, it is likely to contain the key information for answering the question. Please follow the answer format: “Answer: {answer_text}”
Question:
<image>
Non-Clue
Please answer the question regarding a visual feature of an organism (animal, plant, etc.). You will be provided with a image regarding that organism. Please follow the answer format: “Answer: {answer_text}”
Question:
<image>
Multiple Images
Please answer the question regarding a visual feature of an organism (animal, plant, etc.). You will be provided with several images, all of them are regarding that organism, but not all images contain the key information for answering the question. Only one to few images, or even none of them can be used to answer. Please follow the answer format: “Answer: {answer_text}”
Question:
<image> <image> <image>…
Table 7: Prompt Template for Main Question Answering Experiments of Different Settings.
Prompt Template For Automatic Evaluation
Please evaluate the answer to a question, score from 0 to 1. The reference answer is provided, and the reference is usually short phrases or a single keyword. If the student answer is containing the keywords or similar expressions (including similar color), without any additional guessed information, it is full correct. If the student answer have missed some important part in the reference answer, please assign partial score. Usually, when there are 2 key features and only 1 is being answered, assign 0.5 score; if there are more than 2 key features, adjust partial score by ratio of correctly answered key feature. The reference answer can be in the form of a Python list, in this case, any one of the list item is correct.
If student answer contain irrelevant information not related to question, mark it with “Redundant”, but it does not affect score if related part are correct. (e.g. Question: what shape is leave of Sanguinaria canadensis, Student Answer: shape is xxx, color is yyy, this is Redundant answer)
If student answer contain features not listed in reference answer, mark it with “Likely Hallucination” and deduct 0.5 score. (e.g., Reference Answer: black and white. Student Answer: black white, with yellow dots, “yellow dots” is not mentioned in reference)
Separate the remarks with score using “|”, that is, use the syntax of: “Score: score | Likely Hallucination”, “Score: {score}”, “Score: {score} | Likely Hallucination | Redundant”, etc. If any explanation on why giving the score is needed, do not start a new line and append after remark with brackets, e.g. “Score: {score} | Redundant | (Explanation: abc)”.
Following are few examples:
Question: Is there any specific color marking around the eyes of a semipalmated plover (scientific name: Charadrius semipalmatus)?
Reference Answer: black eye-round feather, white stripe above eyes, sometimes connected to the white forehead
Student Answer: Yes, the bird has a distinctive black line that runs through the eye, which is a key identifying feature.
Score: 0 | Likely Hallucination
Student Answer: They have a black vertical band in front of the eye, a white band above the eye, and a single black band that wraps partially around the eye, creating a partial “mask” appearance.
Score: 1
Student Answer: Yes, the semipalmated plover has a distinctive black/dark ring around its eye, surrounded by a bright white ring or patch
Score: 0.5 | Likely Hallucination (Explanation: not white ring, but only a line above the eye)
Question: …
More ICL Examples omitted
Now, Score the following question:
Table 8: Prompt Template For Automatic Evaluation. Note that for evaluation, no images is provided, and it becomes pure-text open-ended QA evaluation. Texts in italic and purple are not part of prompt.
Appendix E Detailed Results
This section contains additional experimental results. We also include ROUGE-1 recall scores [^9] Implemented with https://pypi.org/project/rouge/, version 1.0.1 for readers who are interested. However, because ROUGE relies on strict word matching, the results for open-ended QA can vary significantly due to the different generation styles across models.
Model Qwen2VL Llama3.2-V GPT-4o
Same Irrel Same Irrel Same Irrel
k=3 40.72 36.27 37.00 36.75 43.90 39.79
5 39.68 37.53 35.92 32.16 45.90 45.03
7 39.36 35.66 36.53 29.55 45.31 40.36
10 39.16 33.30 34.51 27.80 46.14 45.29
15 39.40 35.56 34.59 28.19 46.54 45.64
20 37.73 32.12 - - 45.62 44.50
Table 9: Impact from using highly irrelevant distractor in 1-in-k setting. “Same” denotes hard-negatives from same species, “Irrel” denotes irrelevant images from other species.
Setting Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
Zeroshot 27.57 32.04 26.17 30.21 35.80 47.21 42.78 42.43
Oracle 33.67 44.82 38.38 35.96 43.89 38.45 19.6 33.67
34.26 43.63 37.05 37.25 42.82 39.24 18.49 32.47
30.88 43.82 39.37 39.44 37.12 37.25 18.22 27.69
32.67 40.64 39.97 39.04 45.08 36.25 19.55 33.47
33.07 43.03 38.11 37.25 38.65 37.05 16.94 33.67
AVG 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.19
STD 1.285 1.563 1.136 1.433 3.449 1.188 1.096 2.566
Non-clue 26.69 33.27 31.21 33.27 30.28 14.74 4.46 8.17
27.49 34.26 34.06 35.86 34.66 17.73 7.41 8.76
30.28 33.86 32.07 37.05 29.88 17.93 6.71 10.16
27.09 33.47 29.68 34.66 32.53 16.53 7.07 8.96
27.29 31.47 30.61 33.86 30.81 16.14 5.59 8.57
AVG 27.77 33.27 31.53 34.94 31.63 16.61 6.25 8.92
STD 1.435 1.073 1.663 1.527 1.972 1.296 1.211 0.750
Table 10: Baselines detailed results
1 in K Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
k=3 32.27 40.44 37.58 37.45 39.04 41.63 42.94 32.27
32.27 39.64 36.52 35.26 34.79 44.22 49.09 32.67
34.06 40.24 37.71 37.85 34.53 44.02 47.43 35.26
33.86 41.63 36.12 39.64 38.84 43.03 41.99 30.68
32.67 41.63 35.65 35.86 37.78 46.61 46.75 33.47
AVG 33.03 40.72 36.72 37.21 37.00 43.90 45.64 32.87
STD 0.871 0.885 0.903 1.732 2.187 1.828 3.039 1.678
k=5 33.86 39.84 34.39 40.93 37.78 46.81 47.8 37.25
33.20 39.64 34.46 38.05 33.20 46.02 46.54 34.86
34.46 40.24 35.45 44.2 35.06 47.01 46.61 36.06
33.67 39.64 35.99 36.85 37.12 44.42 43.61 35.26
34.66 39.04 35.79 38.45 36.45 45.22 47.18 36.45
AVG 33.97 39.68 35.22 39.70 35.92 45.90 46.35 35.98
STD 0.594 0.434 0.748 2.923 1.825 1.087 1.613 0.950
k=7 31.08 38.65 35.45 34.66 37.91 43.03 46.25 37.4
32.27 39.44 34.13 37.05 35.39 44.02 44.97 35.88
32.27 38.84 36.32 33.47 34.59 45.29 44.71 36.83
36.85 39.44 35.06 36.65 35.85 46.22 45.66 35.88
34.26 40.44 33.06 36.46 38.91 48.01 45.71 37.02
AVG 33.35 39.36 34.80 35.66 36.53 45.31 45.46 36.60
STD 2.267 0.699 1.253 1.530 1.809 1.935 0.618 0.690
k=10 32.27 39.04 35.72 37.00 33.53 45.62 46.73 34.15
33.07 39.64 37.05 37.10 34.45 44.82 48.79 37.65
32.87 39.84 35.12 37.70 36.39 46.41 49.66 38.25
31.47 38.45 35.78 36.23 33.20 47.61 43.54 39.24
31.67 38.84 34.39 34.82 34.99 46.22 45.76 38.84
AVG 32.27 39.16 35.61 36.57 34.51 46.14 46.90 37.63
STD 0.707 0.573 0.980 1.109 1.269 1.031 2.440 2.034
k=15 31.27 40.44 33.86 42.01 31.94 45.42 47.6 34.15
33.86 39.24 35.99 36.55 34.92 48.01 49.63 37.79
30.08 40.24 34.66 36.38 35.52 47.81 42.42 38.74
32.47 38.84 36.65 36.38 36.58 46.02 47.6 39.5
33.47 38.25 33.79 35.27 34.00 45.42 45.66 39.12
AVG 32.23 39.40 34.99 37.32 34.59 46.54 46.58 37.86
STD 1.565 0.929 1.283 2.672 1.754 1.280 2.717 2.169
k=20 33.47 38.05 33.93 - - 44.82 51.15 35.67
29.48 39.44 35.32 - - 44.62 47.99 34.76
33.47 36.25 31.94 - - 45.22 47.13 36.39
32.47 37.65 34.00 - - 47.21 48.57 33.47
33.07 37.25 35.19 - - 46.22 53.01 36.85
AVG 32.39 37.73 34.08 - - 45.62 49.57 35.43
STD 1.679 1.168 1.358 - - 1.083 2.438 1.350
Table 11: 1-in-K detailed results
Setting Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
Zeroshot 14.66 14.87 17.03 20.44 14.05 11.96 12.58 9.65
Oracle 15.26 26.75 17.16 25.83 6.72 16.48 16.35 24.59
15.85 26.63 19.01 25.96 8.1 15.9 14.1 23.26
15.22 25.86 18.72 29.25 7.19 18.44 14.77 23.63
15.79 25.08 19.46 28.42 10.39 20.07 14.58 23.83
15.27 26.44 17.06 25.86 7.88 20.36 11.47 26.57
AVG 15.48 26.15 18.28 27.06 8.06 18.25 14.25 24.38
STD 0.313 0.690 1.102 1.644 1.416 2.028 1.771 1.319
Non-clue 9.49 21.59 8.07 23.25 3.95 9.23 9.04 17.11
9.27 21.13 9.24 23.94 5.4 7.86 10.03 17.18
8.89 20.96 10.78 24.4 2.87 10.06 9.21 17.42
9.85 21.23 8.69 24.63 5 6.62 11.4 15.26
9.27 20.86 9.05 24.18 4.05 9.9 10.14 16.09
AVG 9.35 21.15 9.17 24.08 4.25 8.73 9.96 16.61
STD 0.351 0.283 1.006 0.530 0.990 1.466 0.938 0.912
Table 12: Baselines rouge scores for reference. Please note that rouge score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.
1 in K Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
k=3 10.49 23.80 8.79 24.31 6.92 26.23 16.14 23.27
10.30 25.47 9.65 24.12 6.64 20.81 18.71 22.65
10.57 23.89 8.22 25.83 5.55 24.87 19.38 22.69
9.89 24.72 8.78 25.27 7.97 23.78 17.91 22.74
10.29 25.80 9.34 25.79 6.76 24.18 19.19 21.31
AVG 10.31 24.74 8.96 25.06 6.77 23.97 18.27 22.53
STD 0.263 0.903 0.554 0.809 0.861 1.999 1.317 0.728
k=5 12.18 25.82 7.77 25.82 8.71 22.44 16.94 22.37
9.87 23.39 8.22 25.45 4.79 21.04 16.36 23.2
10.50 23.54 8.65 26.76 5.55 25.28 16.95 23.62
11.76 22.36 7.72 22.11 7.65 17.79 16.17 23.17
10.30 24.67 8.62 24.56 5.12 28.04 17.84 24.35
AVG 10.92 23.72 8.20 24.94 6.36 22.92 16.85 23.34
STD 0.995 0.967 0.446 1.768 1.721 3.933 0.652 0.722
k=7 9.28 23.26 8.63 21.65 7.57 22.62 15.67 24.38
10.72 24.28 8.17 24.61 5.66 25.55 20.08 21.73
11.45 24.74 8.38 21.51 7.14 20.67 16.35 24
12.10 24.81 8.22 24.40 7.63 25.97 17.62 22.79
12.48 24.32 7.67 23.57 7.01 24.4 18.82 23.07
AVG 11.21 24.28 8.21 23.15 7.00 23.84 17.71 23.19
STD 1.268 0.619 0.353 1.484 0.796 2.197 1.794 1.046
k=10 11.85 23.90 7.65 25.71 5.81 23.86 18.92 23.03
12.15 24.00 8.81 25.50 6.93 24.44 17.5 21.05
11.83 23.69 8.19 24.96 6.15 23.38 18.64 21.52
11.79 24.45 8.77 25.28 6.62 24.34 17.57 23.31
11.06 24.63 8.06 26.36 6.78 26.39 19.17 21.71
AVG 11.74 24.13 8.30 25.56 6.46 24.48 18.36 22.12
STD 0.404 0.392 0.493 0.525 0.466 1.147 0.776 0.990
k=15 11.42 23.21 8.31 24.68 7.44 21.5 19.12 20.55
11.52 24.76 8.93 25.77 8.41 25.05 19.79 21.93
10.14 24.49 7.84 26.46 7.07 23 18.47 24.1
11.77 23.49 8.11 23.52 7.89 28.41 20.25 24.11
11.77 23.34 8.52 26.43 6.96 21.25 20.24 23
AVG 11.32 23.86 8.34 25.37 7.55 23.84 19.57 22.74
STD 0.680 0.714 0.414 1.262 0.601 2.968 0.770 1.520
k=20 13.09 22.95 8.02 - - 24.21 18.57 20.79
11.51 23.58 9.40 - - 21.31 16.66 20.56
11.76 23.24 8.30 - - 22.27 17.95 23.95
11.85 23.94 8.20 - - 21.99 19.25 20.61
11.95 22.80 9.00 - - 25.61 20.19 23.31
AVG 12.03 23.30 8.58 - - 23.08 18.52 21.84
STD 0.614 0.465 0.588 - - 1.779 1.333 1.648
Table 13: 1-in-K rouge scores for reference. Please note that rouge score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.

### v2

@@
 Abstract
-Retrieval-Augmented Generation (RAG) is a popular approach for enhancing Large Language Models (LLMs) by addressing their limitations in verifying facts and answering knowledge-intensive questions.
-As the research in LLM extends their capability to handle input modality other than text, e.g.
-image, several multimodal RAG benchmarks are proposed.
-Nonetheless, they mainly use textual knowledge bases as the primary source of evidences for augmentation.
-There still lack benchmarks designed to evaluate images as augmentation in RAG systems and how they leverage visual knowledge.
-We propose Visual-RAG, a novel Question Answering benchmark that emphasizes visual knowledge intensive questions.
-Unlike prior works relying on text-based evidence, Visual-RAG necessitates text-to-image retrieval and integration of relevant clue images to extract visual knowledge as evidence.
-With Visual-RAG, we evaluate 5 open-sourced and 3 proprietary Multimodal LLMs (MLLMs), revealing that images can serve as good evidence in RAG; however, even the SoTA models struggle with effectively extracting and utilizing visual knowledge [^1] Benchmark and evaluation codes will be available at https://github.com/LuciusLan/Visual-RAG..
-Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries
-Yin Wu1, Quanyu Long1, Jing Li2, Jianfei Yu3, Wenya Wang1 1 Nanyang Technological University, 2 Harbin Institute of Technology (Shenzhen), 3 Nanjing University of Science and Technology wuyi0023@e.ntu.edu.sg
+Retrieval‑augmented generation (RAG) is a paradigm that augments large language models (LLMs) with external knowledge to tackle knowledge‑intensive question answering.
+While several benchmarks evaluate Multimodal LLMs (MLLMs) under Multimodal RAG settings, they predominantly retrieve from textual corpora and do not explicitly assess how models exploit visual evidence during generation.
+Consequently, there still lacks benchmark that isolates and measures the contribution of retrieved images in RAG.
+We introduce Visual-RAG, a question‑answering benchmark that targets visually grounded, knowledge‑intensive questions.
+Unlike prior work, Visual-RAG requires text‑to‑image retrieval and the integration of retrieved clue images to extract visual knowledge necessary for answer generation.
+With Visual-RAG, we evaluate 5 open-source and 3 proprietary MLLMs and demonstrate that images provide strong evidence in augmented generation; however, even state‑of‑the‑art models struggle to effectively extract and utilize visual knowledge.
+Our results highlight the need for improved visual retrieval, grounding, and attribution in multimodal RAG systems.
+[^1] Benchmark available at: github.com/visual-rag/visual-rag
 1 Introduction
-Figure 1: Comparison on query format and QA pipeline between InfoSeek-style and our Visual-RAG.
-InfoSeek-style: The entity in the query image must first be recognized to extract textual evidence, which is then used independently to answer the question.
-Visual-RAG: The query is text-only, but direct answers are rarely available in textual form.
-The model must retrieve relevant images and interpret them to generate an answer.
-Benchmark Query Modality Knowledge Modality Retrieval Modality InfoSeek (Chen et al., 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] E-VQA (Mensink et al., 2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] WebQA (Chang et al., 2022) [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] [Uncaptioned image] MRAG-Bench (Hu et al., 2024) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Visual-RAG (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
-Table 1: Comparing modalities of knowledge intensive VQA benchmarks.
+Figure 1: InfoSeek style: Textual-Knowledge-Centric RAG (left), MRAG-Bench style: Retrieving common images for entity recognition and entity bridged prior knowledge recalling (middle) v.s.
+Ours: Visual Evidence-Centric Retrieval and Visual Knowledge-Intensive Reasoning (right).
+MRAG-Bench (Hu et al.
+2025) formulates its task by presenting uncommon images that are challenging for entity recognition, retrieving more common images to support the MLLMs in bridging object identification.
+Typically, the actual answers (e.g., identifying the engine used in a specific car) is rarely explicitly represented visually in the retrieved images, making the answer dependent heavily on the model’s internal prior knowledge.
+In contrast, our Visual-RAG benchmark utilizes text-only queries with a common identifier (e.g., name of an organism), and requires retrieval of specialized, less common images containing explicit answer within their visual contents.
+Consequently, Visual-RAG makes cross-modal retrieval more challenging and demands deeper visually grounded knowledge acquisition, thus facilitating a more precise visual-centric evaluation of MLLMs’ capability to extract detailed visual information for knowledge-intensive reasoning.
+Retrieval‑Augmented Generation (RAG) augments large language models (LLMs) with external knowledge, substantially improving performance on knowledge‑intensive queries (Lewis et al.
+2020; Karpukhin et al.
+2020; Izacard and Grave 2021).
+While the application of RAG is well-acknowledged for text-only knowledge-intensive tasks, many real‑world scenarios require multimodal knowledge from visual cues, such as diagrams or maps, highlighting the limitations of relying solely on text retrieval.
+Several benchmarks have been proposed to evaluate Multimodal Large Language Models (MLLMs) under multimodal RAG settings, targeting complex, knowledge‑intensive questions.
+Notably, InfoSeek (Chen et al.
+2023) and Encyclopedic-VQA (Mensink et al.
+2023) pair mixed‑modal queries directly with images.
+Yet, these benchmarks are fundamentally Textual‑Knowledge‑Centric: answering typically involves first recognizing entities in the query image and then retrieving textual evidence (e.g., Wikipedia passages) to generate the answer, as demonstrated in Figure 1 (left).
+As a result, although images are present in query, MLLMs rarely need to extract visual knowledge from them during answer generation, and current multimodal RAG benchmarks predominantly assess textual rather than visual knowledge exploitation.
+Recent work MRAG-Bench (Hu et al.
+2025) partially addresses this gap by explicitly requiring image retrieval.
+As shown in Figure 1 (middle), MRAG-Bench features challenging images as queries (e.g., damaged, partially occluded or incomplete entities), enforcing models to retrieve more common images of the entity to assist entity recognition.
+However, final answers (e.g., the engine used in a specific car model) largely rely on parametric model knowledge rather than explicit visual evidence contained in the retrieved images.
+Consequently, solving the task still depends on internal priors instead of information extracted from the pictures, leaving the core challenge of visually grounded RAG unaddressed.
+We therefore introduce Visual-RAG, a visual knowledge‑intensive QA benchmark that explicitly requires text‑to‑image retrieval and visually grounded answer generation.
+Crucially, the retrieved image itself serves as evidence, which aligns with real-world use case: once a clue image containing the relevant visual feature is surfaced, one can answer the visual knowledge-intensive question directly, without treating the image as a pivot to trigger subsequent recalling of prior knowledge.
+Accordingly, Visual-RAG emphasizes Visual Evidence‑Centric Retrieval and mandates Visual Knowledge‑Intensive Reasoning, challenging models to extract and ground the information contained in clue images, and to discriminate it from visually similar hard-negative distractors.
+Key features of Visual-RAG include:
+  1. 1.
+    Fine‑grained and less common visual evidence. Only 6.12% of corpus images are clue images containing the necessary visual evidence, substantially making cross‑modal retrieval challenging.
+  2. 2.
+    Hard negatives at scale.
+Most non‑clue images are hard negatives, which depict the same entities and are visually similar, stress‑testing both retrieval models’ robustness and downstream MLLMs’ capabilities in visual knowledge extraction.
+  3. 3.
+    Explicit visual answerability. Evidences of answers are encoded directly in the visual content, enabling evaluation of whether MLLMs truly leverage visual information.
+To construct Visual-RAG, we initially generate candidate queries using LLMs, followed by rigorous refinement through human filtering and rewriting processes.
+Subsequently, we employ both open-source MLLMs and human annotators for additional filtering, ensuring the final queries reflect genuine human curiosity and practical relevance.
+Benchmark Query Modality Knowledge Modality Retrieval Modality
+InfoSeek (Chen et al.
+2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+E-VQA (Mensink et al.
+2023) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+WebQA (Chang et al. 2022) [Uncaptioned image] [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+MRAG-Bench (Hu et al. 2025) [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+Visual-RAG (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image]
+Table 1: Comparing modalities of knowledge-intensive VQA benchmarks.
 The link icon [Uncaptioned image] denotes that the two objects are paired, retrieving one will automatically link to the other.
 For InfoSeek and E-VQA, Yan and Xie (2024) demonstrated that by image-to-image retrieval on Wikipedia images, the Wikipedia articles containing those images are also retrieved at a significantly higher recall comparing to image-to-text retrieval.
-Similarly for WebQA, a question-to-caption (text-to-text) retrieval achieves better performance than text-to-image retrieval, as shown in Table 5 in Appendix.
+Similarly for WebQA, a question-to-caption (text-to-text) retrieval achieves better performance than text-to-image retrieval, as shown in Appendix A.
 Our Visual-RAG enforces text-to-image retrieval for visual knowledge.
-We have observed rapid advancement of Large Language Models (LLMs) in recent years.
-However, they are prone to generating hallucinated responses Ji et al.
-(2023).
-Retrieval-Augmented Generation (RAG) has emerged as a widely adopted approach to mitigate this challenge (Lewis et al., 2020; Karpukhin et al., 2020; Izacard and Grave, 2021).
-By retrieving and integrating relevant textual documents from external knowledge sources, it enables LLMs to answer knowledge-intensive questions more accurately.
-Yet, many real-world scenarios demand multimodal knowledge (e.g., visual cues or geographical maps), suggesting that text-only retrieval may be insufficient.
-Building on the success of text-only RAG, various evaluation benchmarks have been proposed to assess the capability of Multimodal LLMs (MLLMs, a.k.a.
-Large Vision Language Models LVLM) in using multimodal RAG to address complex, multimodal knowledge intensive questions, including InfoSeek (Chen et al., 2023), Encyclopedic-VQA (Mensink et al., 2023), etc.
-They feature mixed-modal queries where the question is asked based on a paired image.
-However, existing datasets reveal several limitations.
-Firstly, although images are included in the query, MLLMs do not utilize the visual knowledge from these images when generating answers.
-Secondly, questions in these datasets typically do not require reasoning over knowledge extracted from images.
-Figure 1 gives an example on InfoSeek-styled query.
-To answer the query, a model needs to first recognize the entity in query image, then locate the Wikipedia article and augment relevant passage to LLMs to generate an answer.
-Query image information is only required in the entity recognition step.
-These existing multimodal benchmarks are textual knowledge intensive, highlighting the need for a new benchmark for visual knowledge intensive QA.
-To this end, We present Visual-RAG, a visual knowledge intensive QA benchmark designed to evaluate text-to-image retrieval and answer generation augmented by retrieved images.
-Key features of Visual-RAG are:
-  1. 1.
-    The entities in queries are obscure, with their descriptions being brief, rarely documented in standard references (e.g., Wikipedia)
-  2. 2.
-    The visual features queried are fine-grained, with only 5.35% of the images in the corpus containing the specific visual evidence needed to answer the questions, posing a challenging text-to-image retrieval task.
-  3. 3.
-    The majority of non-clue images are hard-negatives which depict the same entities and are visually similar. This subtle distinction further increases the difficulty of cross-modal retrieval.
-To construct Visual-RAG, we begin with LLM-generated candidate queries, which are then refined through human filtering and rewriting.
-Next, we employ open-sourced MLLM and human annotators to perform another filtering step to ensure an appropriate level of difficulty for the queries.
-Detailed procedures for the entire benchmark construction can be found in Section 4.
-With Visual-RAG, we comprehensively evaluate 8 mainstream MLLMs, including 5 open-sourced models and 3 proprietary ones.
-Firstly, we demonstrate that images can serve as powerful evidence for augmented generation through baseline experiments; secondly, we experiment with various RAG settings, testing the ability of MLLMs digesting images as augmentation, and analyse model behaviour on identifying clue from irrelevant images.
-Our benchmark sheds light on real-world scenarios for evaluating MLLMs, when users raise natural language queries for knowledge-intensive questions where textual evidence is scarce but visual evidence possibly exists, posing the challenging task of retrieving and utilizing relevant images in the wild.
+With Visual-RAG, we comprehensively evaluate eight popular MLLMs, including five open-source and three proprietary models.
+Our baseline results confirm that images can serve as strong evidence for augmented generation.
+We also examine multiple RAG configurations to assess how effectively MLLMs ingest image augmentation, and to analyze their ability to discriminate informative visual clues from irrelevant distractors.
 The key findings from our experiments are summarized as follows:
-  • Cross-modal retrieval is challenging.
-The commonly used small cross-modal retriever (CLIP) faces significant challenges addressing difficult queries requiring the identification of fine-grained visual features.
-  • The amount of evidence affects retrieval-augmented generation.
-The evaluated MLLMs demonstrate the ability to extract visual knowledge to answer questions.
-However, a single ground-truth clue image is insufficient for the models to achieve optimal performance.
-  • Negatives can be helpful.
-Providing multiple images as input, including both clue images and non-clue images, could enhance performance for proprietary models.
-This improvement likely stems from the models’ ability to contrast clue and non-clue images, enabling them to better process the relationship between the question and the images.
-  • Inverted dynamics of open-sourced and proprietary models.
-Open-source models excel with a single clue image but fail to differentiate the clue when irrelevant images are also introduced; while proprietary models are initially conservative with a single clue but ultimately achieve better performance through multi-image augmentation.
-This reveals open-source models are weaker at handling multiple images, compared to proprietary models.
+  • Visual evidence-centric retrieval is challenging.
+The commonly used cross-modal retrievers face significant challenge in addressing difficult queries requiring the identification of fine-grained visual evidence.
+  • Visual clues can serve as effective evidence for Retrieval-Augmented Generation.
+The evaluated MLLMs demonstrate the ability to extract visual knowledge to answer questions when presented with single ground-truth clue image.
+  • Realistic visual RAG remains an onerous task.
+With top-$k$ RAG, performance of most models lags behind the ground-truth clue upper bound.
+Proprietary models reach the upper bound only when $k$ close to 20, revealing limited ability to efficiently exploit visual clues, partially constrained by the bottleneck of retriever.
+  • Inverted dynamics of open-source and proprietary models.
+When a guaranteed clue is mixed with non-clue images, open-source MLLMs leverage it well at lower $k$, but their accuracy decays as including more distractors.
+In contrast, proprietary MLLMs retain robustness in identifying visual evidence from distractors at higher $k$, yet overlook the clue when only a few images are provided.
 2 Related Work
-While numerous benchmarks evaluate the capabilities of MLLMs, we focus here on those designed for knowledge-intensive visual question answering (VQA), where RAG is the primary solution paradigm.
-Table 1 compares the modalities covered by commonly used benchmarks.
-A more detailed discussion of how Visual-RAG differs from prior work is provided in Appendix A.
-Textual Knowledge Centric VQA Benchmarks
-OK-VQA (Marino et al., 2019) and its augmented version A-OKVQA (Schwenk et al., 2022) focus on commonsense or world knowledge, though they do not strongly depend on external knowledge retrieval.
-ViQuAE (Lerner et al., 2022), InfoSeek (Chen et al., 2023) and Encyclopedic-VQA (E-VQA, Mensink et al., 2023) address “encyclopedic” questions, paired with query images of specific entities, utilizing information retrieved from external textual sources (e.g., Wikipedia) to answer.
-Visual Knowledge Centric QA Benchmarks
-While the above benchmarks all position themselves as knowledge-intensive VQA tasks, it is important to note that the required knowledge remains primarily textual, with images functioning mostly as entity anchors.
-In contrast, to our knowledge, only two related works truly require image retrieval as evidence for answering questions.
-WebQA (Chang et al., 2022) is an earlier effort that requires retrieving images for visual knowledge.
-However, it heavily relies on captions because the images themselves usually lack indicators of which are clue images.
-A concurrent work, MRAG-Bench (Hu et al., 2024) shares a similar motivation in using image retrieval for question answering.
-However, it primarily aims to recognize an entity when the instance in query image is incomplete or partially obscured, relying on retrieved images that present a complete view of the same entity.
-This entity-recognition objective differs from our focus on answering visual-feature-oriented factual queries.
-As queries in MRAG-Bench include image, the model can use the “shortcut” of matching a visually similar retrieved candidate, bypassing the need for deeper text-to-image reasoning.
+Numerous benchmarks have been developed to evaluate the capabilities of MLLMs.
+We focus on those specifically designed for knowledge-intensive visual question answering (VQA), particularly emphasizing retrieval-augmented generation (RAG) paradigms.
+Table 1 provides an overview and comparison of modalities across widely adopted benchmarks.
+Textual Knowledge Centric VQA Benchmarks.
+Benchmarks such as OK-VQA (Marino et al.
+2019) and its augmented variant, A-OKVQA (Schwenk et al.
+2022), predominantly emphasize commonsense and world knowledge but exhibit limited dependence on external knowledge retrieval mechanisms.
+Other datasets like ViQuAE (Lerner et al.
+2022), InfoSeek (Chen et al.
+2023) and Encyclopedic-VQA (E-VQA, Mensink et al.
+2023) address encyclopedic questions paired with images of specific entities.
+However, the retrieved information employed in answering is primarily textual (e.g., passages from Wikipedia), with the visual modality mainly serving as an entity anchor rather than as a source of knowledge.
+Visual Knowledge Centric QA Benchmarks.
+To date, only a limited number of benchmarks explicitly require visual knowledge retrieval as evidence in answering questions.
+WebQA (Chang et al.
+2022) is an earlier effort that necessitates retrieving images for visual knowledge but relies heavily on textual captions, as images in the benchmark often lack explicit indicators for being clue.
+More recently, MRAG-Bench (Hu et al.
+2025) similarly emphasizes image retrieval.
+However, MRAG-Bench provides queries in the form of challenging images depicting partially obscured or incomplete entities, aiming primarily at entity recognition through matching visually similar images.
+The retrieved images predominantly help in identifying entities, with answers typically dependent on the model’s prior knowledge rather than directly on visual evidence within these images.
+Unlike MRAG-Bench, our benchmark, Visual-RAG , explicitly constructs textual queries and requires retrieving specialized images that explicitly contain visual evidence.
+Consequently, our benchmark evaluates deeper visual knowledge extraction capabilities for MLLMs, emphasizing visual evidence-centric reasoning.
+More detailed comparison between Visual-RAG and prior benchmarks is provided in Appendix A.
 3 Benchmark Overview
-We propose Visual-RAG, a text-only Question Answering benchmark for challenging, visual knowledge intensive factual questions.
-It is an evaluation benchmark for visual knowledge centric text-to-image retrieval, as well as a benchmark for evaluating MLLMs’ abilities on extracting visual knowledge as augmentation under retrieval-augmented generation settings.
-Visual Knowledge Intensive Queries
-Our Visual-RAG focuses on text-only queries that probe factual knowledge of visual feature which generally hold true of an entity (e.g.
-“What colour are the stamens of Chamaecrista nictitans”), rather than instance-specific attributes commonly addressed in standard VQA tasks (e.g.
-“What colour is the cup on table”).
-Hence, we specifically select the organism domain which enables specialized queries.
-Unlike existing works, we deliberately exclude images from the queries due to the reason that including an image often reduces the problem to mere entity recognition or unimodal-similarity match, instead of our main objective of cross-modal knowledge extraction.
-Numerous established benchmarks already cover entity recognition, including the OVEN (Hu et al., 2023) dataset which serves as the foundation for the the InfoSeek dataset, and the iNaturalist 2021 (Van Horn et al., 2021) dataset, which we employ in this work.
-Moreover, text-only queries better align with real-world scenarios where users tend to raise questions about an entity without an image.
-Naturally Co-occurring Hard Negative Images
-For each question, multiple clue images might be available, reflecting real-world scenarios.
-Additionally, the dataset incorporates a substantial number of naturally co-occurring hard negative images, i.e., images of the same species that do not exhibit the visual feature in query.
-Notably, all clue images for a given question can be retrieved purely based on visual information, without reliance on captions.
-Dataset Statistics
-The current version of Visual-RAG contains 400 queries, forming an image knowledge base of total 103824 images.
-The average clue image count for each query is 13.89, and the clue images occupy 5.35% of the image knowledge base.
-The distribution of organism categories and query types is shown in Figure 2.
-(a) Distribution of organism categories in Visual-RAG
-(b) Distribution of question categories
-Figure 2: Distribution of organisms and question categories.
+We propose Visual-RAG, a Question Answering (QA) benchmark for visual knowledge–intensive factual queries.
+Visual-RAG evaluates two capabilities under retrieval‑augmented generation (RAG): (i) visual evidence–centric text‑to‑image retrieval and (ii) an MLLM’s ability for visual knowledge-intensive reasoning with retrieved images in retrieval-augmented answer generation.
+Task Definition.
+Let $\mathcal{Q}$ be a set of text queries and $\mathcal{D}$ a corpus of images.
+For each $q\in\mathcal{Q}$ there exists a small set of clue images $\mathcal{I}_{c}\subset\mathcal{D}$ whose pixel content explicitly encodes the visual evidence needed to answer $q$.
+The remaining images $\mathcal{I}_{d}=\mathcal{D}\setminus\mathcal{I}_{c}$ act as distractors; many are hard negatives that depict the same entity or highly similar instances but lack the queried visual feature.
+Given $q$, a system first retrieves $k$ images $\mathcal{I}_{ret;k}\subset\mathcal{D}$ and then generates an answer $\hat{y}$ conditioned on $(q,I_{ret;k})$.
+Crucially, Visual-RAG assumes that once a correct clue image is retrieved, the answer is directly derivable from that image, without requiring subsequent textual lookup.
+Visual Knowledge–Intensive Queries.
+Queries in Visual-RAG are text‑only.
+They probe factual properties of visual features that generally hold for an entity category (e.g., “What color are the hindwings of Papaipema inquaesita?”, as in Figure 1), rather than instance‑specific properties typical of standard VQA (e.g., “What color is the cup on the table?”).
+We focus on the organism domain because it naturally supports fine‑grained, specialized questions for which the necessary evidence is visual and stable at the category level.
+Unlike many existing works, we deliberately exclude image from the query input to avoid reducing the task to entity recognition or image-to-image unimodal similarity matching; the primary objective is cross‑modal knowledge extraction from retrieved images.
+Numerous established benchmarks already cover entity recognition, including the OVEN (Hu et al.
+2023) dataset which serves as the foundation for the InfoSeek dataset, and the iNaturalist 2021 (Van Horn et al.
+2021) dataset, which we employ in this work.
+Moreover, our setting of visual knowledge-intensive queries simulates realistic application scenarios, where referring directly to images for answers is more intuitive and efficient compared to searching through textual documentation for such query.
+Naturally Co-occurring Hard Negative Images.
+Visual-RAG incorporates a large number of naturally co‑occurring hard negatives: images of the same species that do not exhibit the queried feature.
+All clue images are retrievable based on visual content alone; captions or alt‑text are not required or assumed at retrieval time.
+This construction jointly pressures the retriever - to surface the rare, feature‑bearing images, and the answer generator MLLM - to extract and ground the relevant visual evidence.
+Dataset Statistics.
+Our Visual-RAG contains 374 queries, forming an image knowledge base of total 99,017 images.
+On average, each query has 16.19 clue images; across the corpus, clue images constitute 6.12% of all images.
+The distributions of organism categories and query types are shown in Figure 2.
+(a) Organism categories.
+(b) Question categories.
+Figure 2: Distribution of organisms and question categories in Visual-RAG .
 4 Benchmark Construction
 4.1 Data source
-Following Encyclopedic VQA Mensink et al.
-(2023), we adopt iNaturalist 2021 (iNat21, Van Horn et al., 2021) as the cornerstone of our benchmark.
+Following Encyclopedic VQA (Mensink et al.
+2023), we adopt iNaturalist 2021 (iNat21, Van Horn et al.
+2021) as the cornerstone of our benchmark.
 iNat21 contains 2.7M images of organism (animals, plants, fungi, etc.), spanning across 10000 taxonomy species.
 Each species in iNat21 has 200–300 images, originally collected for species-level image classification.
-For organism species selection in query construction, we aim to ensure that the chosen species are relatively underrepresented in textual resources.
-To achieve that, we filter them based on the length of their Wikipedia summary and description sections.
-Well-known species typically have multiple sections (e.g., habitats, behaviour, etc.), indicating extensive documentation.
-We therefore include only those species whose summary and description account for no more than 50% length of the entire Wikipedia article.
-This also helps minimize reliance on prior knowledge: widely recognized species may have detailed descriptions on visual features across LLM pre-training corpora, even if these details are not explicitly mentioned in Wikipedia.
 4.2 Query Collection
 Question generation.
-Our objective is to create queries about distinct visual features of organisms, ensuring that only a small fraction of images within each species display the queried feature.
-Each question must yield a unique answer and generally hold true for every instance of that species, preserving both specificity and accuracy.
-Typically, composing such specialized queries would require domain expertise and a labour-intensive review of the 200–300 images per species.
-Consequently, we employ OpenAI-o1 (OpenAI, 2024) for query generation to replace human experts.
-As creating meaningful questions requires expert knowledge, we provide the Wikipedia summary and description sections for the model to understand the species.
-We explicitly instruct the model NOT to generate questions about properties already covered in these text passages, prompting it to focus on more specialized visual attributes unlikely to appear in standard references.
-The reasoning and self-validating ability of o1 enables it to follow the complex instructions, leading to high-quality and diverse prototype queries.
-The model is tasked with generating both a set of candidate queries and their corresponding visual features.
-A refined human annotator filtering and rewriting process is then applied based on these visual features, to ensure query validity [^2] More details about the query generation prompt and human filtering process can be found in Appendix C, D..
+Our goal is to create text queries about distinct visual features of organisms such that, for a given species, only a small fraction of images exhibit the queried feature.
+Each query must admit a unique answer or answer set that generally holds at the species level (category‑level fact), rather than instance‑specific properties.
+Constructing such queries by hand would require domain expertise and careful inspection of roughly 200–300 images per species.
+To scale this process, we employ OpenAI‑o1 as a query generator.
+For each candidate species, we provide its Wikipedia Summary and Description sections to the model to convey basic context.
+We explicitly instruct the model NOT to generate questions about properties already described in these passages, pushing it toward specialized visual attributes that are unlikely to appear in standard references.
+The model outputs (i) a set of candidate questions and (ii) their associated visual feature descriptors (e.g., “hindwing color”, “presence of mottled pattern”, etc.).
+To reduce the chance that answers can be trivially recovered from text corpora, we bias species selection toward those that are relatively under‑represented in Wikipedia.
+A subsequent refined human filtering and rewriting process (Appendix C, D) validates the linguistic form of each query and the feature descriptors.
 Question Filtering and Image Annotation.
-The question generation process cannot guarantee compliance with the previously mentioned requirement, i.e., clue images containing the answer should remain a minority in the set; also, query should not be so difficult that no clue image exists in the corpus.
-We designate a threshold clue image rate, that at most 20% of images in a species level sub-corpus can be clue.
-Firstly, we employ an open-sourced MLLM [^3] Llama-3.2-11B-Vision-Instruct was used.
-Prompt template can be found in Appendix D as a coarse-level filter.
-For each query, the model iterates over the sub-corpus one image at a time, binary labelling the images by whether the referenced feature is present.
-These labels are used to roughly filter out queries that exceed the 20% clue-image threshold.
-Our rationale is that if an unusually large fraction of images exhibit a visual feature, it is likely prominent for the MLLM to detect and generate labels that are reliable enough.
-Secondly, we recruit volunteer university students as annotators to label the images.
-These annotators identify whether each image provides evidence for the question and thereby confirm whether (a) at least one valid clue image exists, excluding overly difficult queries having no clue image and (b) fewer than 20% of the images are clue, excluding remaining “easy” queries not recognized by the coarse-filter.
-At the same time, the annotators provide the ground-truth answers for each query.
-This two-stage process ensures our final dataset meets both the clue-image minority requirement and the fundamental criterion that each query should be indeed answerable by evidence clue image within the corpus.
-In the current version of the benchmark, we initially select around 1000 questions that passed the coarse filter.
-After human annotation, 400 of these questions satisfy the clue-image rate threshold and are confirmed answerable with at least one valid clue image.
-5 Experimental Setup
-5.1 Retrieval Evaluation
-For retrieval, we evaluate the performance with Normalized Discounted Cumulative Gain (NDCG) and hit rate – the probability that at least one clue image is found among the top-$k$ retrieved samples (NDCG@k and hit@k).
+Our curation process enforces two constraints: (C1) clue images (images that clearly display the queried feature with the required view/pose) must be a minority within the species‑specific sub‑corpus; and (C2) each query must be answerable by at least one clue image.
+Formally, for a query $q$ over species $s(q)$ with image set $\mathcal{D}_{s(q)}$, let $\mathcal{I}_{c}\subset\mathcal{D}_{s(q)}$ denote the clue set and define the clue rate $\rho_{q}\triangleq\frac{|\mathcal{I}_{c}|}{|\mathcal{D}_{s(q)}|}$.
+We retain only queries within threshold $0<\rho_{q}\leq 0.25$.
+Stage 1: MLLM‑based coarse filter.
+We use an open‑source MLLM [^2] Llama-3.2-11B-Vision-Instruct.
+Prompt template can be found in Appendix D.
+as an image‑level presence / absence classifier for the referenced feature.
+For each query $q$, the model iterates over $\mathcal{D}_{s(q)}$ one image at a time and predicts the binary label indicating whether the feature is visually present.
+From these labels we obtain an estimate clue rate $\hat{\rho}_{q}$ and discard queries with $\hat{\rho}_{q}>0.25$.
+The premise is that when a queried feature is unusually common, it is sufficiently prominent for the MLLM to detect reliably, making the coarse estimate adequate for pre‑screening.
+Stage 2: Human annotation and answer adjudication.
+We recruit volunteer university students as annotators to (a) label images by whether the queried feature is visibly present and (b) provide the ground‑truth answer for the query $q$.
+This step confirms that (i) at least one valid clue image exists and (ii) $\rho_{q}\leq 0.25$, thereby removing queries that slipped past the coarse filter.
+We take the human labels as ground truth for $\mathcal{I}_{c}$ and compute the final $\rho_{q}$ accordingly.
+Approximately 1,0001{,}000 questions pass the coarse MLLM filter.
+After human verification, 374 questions satisfy the clue‑rate criterion ($0<\rho_{q}\leq 0.25$) and are confirmed answerable by at least one clue image.
+5 Experiments
+5.1 Experimental Setups
+Retrieval Evaluation.
+For retrieval, we evaluate the performance with Recall score, Normalized Discounted Cumulative Gain (NDCG), Hit rate and Hit Count.
+Recall@$k$ measures the fraction of all ground-truth clue images that appear within the top-$k$ results, reflecting overall coverage.
+NDCG@$k$ rewards correct images that appear higher in the ranking by discounting relevance logarithmically with position, thus capturing ranking quality.
+Hit@$k$ is a binary success indicator for probability of at least one clue is retrieved in the top-$k$ images, while Hit Count@$k$ tallies how many clue images are found in that cut-off, offering an intuitive sense of how much useful evidence the retriever supplies.
 Due to the challenge of complex domain-specific queries, it is significantly difficult for the retrievers to search through the entire iNat21 corpus (2.7M images) while the query is about a sophisticated feature for a single species.
-To enhance retrieval effectiveness and emphasize the more fine-grained and distinguishing features of a species, we narrow the search space by limiting the retrieval corpus to a set of 200–300 images corresponding to the species for each query.
-In all following experiments, we adopt the commonly used retriever clip-vit-large-patch14-336 (Radford et al., 2021) for image retrieval.
-The search indexes are constructed with the FAISS library (Douze et al., 2024) using flat inner product index.
-5.2 Generation Evaluation
+To enhance retrieval effectiveness and emphasize the more fine-grained and distinguishing features of a species, we narrow the search space by limiting the retrieval corpus to $\mathcal{D}_{s(q)}$, the set of 200–300 images corresponding to the species for each query.
+Dense Retrievers.
+Three dense retriever models are evaluated: (1) The commonly used multimodal retriever CLIP (clip-vit-large-patch14-336, Radford et al.
+2021).
+(2) A recent CLIP variant, BGE (BGE-VL-large, Zhou et al.
+2024.
+(3) MLLM-based retriever, VLM2Vec-qwen2VL-2B (Jiang et al.
+2025).
+The search indexes are constructed with the FAISS library (Douze et al.
+2024) using flat inner product index.
+RAG experiments in following sections adopt images retrieved by CLIP, as CLIP is being adopted as the de-facto retriever in previous multimodal RAG works.
+Generation Evaluation.
 Previous work has relied predominantly on exact-match or recall-based metrics to assess answer correctness.
 However, such methods are prone to overlooking partial hallucinations blended with ground-truth answers.
-For example, if the true answer is “black and white” but the model predicts “black and white with yellow dots,” the spurious mention of “yellow dots” remains undetected under exact-match criteria.
-To address this limitation, we employ GPT-4o as a more nuanced evaluator.
+For example, if the true answer is “black and white” but the model predicts “black and white with yellow dots”, the spurious mention of “yellow dots” remains undetected under exact-match criteria.
+To address this limitation, we adopt LLM-as-Judge, employing OpenAI o4-mini as a more nuanced evaluator.
+We retain all annotator-provided answers as valid ground-truth candidates, and specify in instruction prompt to consider alternative expressions of features.
 Detailed evaluation prompt template can be found in Appendix D.
@@
 (2024), have demonstrated that general-purpose LLMs can reliably judge open-ended QA responses.
-We also manually validate the automatic evaluation results on a sample of 100 predicted answers, achieving a 94% accuracy rate.
-In addition, we report the ROUGE score (Lin, 2004), a metric commonly used in summarization evaluation, as a relaxed version of exact match score.
+We also manually evaluated 200 sampled predicted answers, and the LLM labels yielded alignment score of $F_{1}=0.92$ with human annotator.
+The confusion matrix can be found in Appendix  E.
+In addition, we report the ROUGE score (Lin 2004), a metric commonly used in summarization evaluation, as a relaxed version of exact match score.
 The ROUGE scores are omitted in result analysis section, and can be found in Appendix E.
-5.3 Multimodal LLMs
-We evaluate 3 advanced proprietary models: GPT-4o (OpenAI, 2025), Gemini-1.5-Pro (Google, 2025), Claude-3.5-Sonnet (Anthropic, 2025); and 5 recent open-sourced models that can process multiple images, of sizes ranging from 4B to 12B: Phi-3.5-vision-instruct (4B, Microsoft, 2024), Qwen2-VL-7B-Instruct (Wang et al., 2024), InternVL-2.5-8B (Chen et al., 2024), Llama-3.2-11B-Vision-Instruct (Meta Llama Team, 2024), and Pixtral-12B (Mistral AI, 2024).
-We use the default setting given by each model to perform question answering.
-6 Results
-6.1 Text-to-Image Retrieval
-Metric @1 @5 @10 @20 @30
-NDCG 26.06 26.46 30.67 38.50 45.80
-Hit Rate 26.06 52.73 64.55 72.73 77.87
-Hit Count 0.26 1.12 2.08 3.60 4.85
-Table 2: Retrieval Results using the sub-corpus of 200-300 images for each query. Even within the small species level corpus, the model struggles with our challenging text-to-image retrieval task.
+Multimodal LLMs.
+We evaluate three flagship proprietary models: GPT-4o (OpenAI 2025), Gemini-2.5-Pro (Google 2025), Claude-Sonnet-4 (Anthropic 2025); [^3] GPT-4o does not internally implement long Chain-of-Though reasoning, while Gemini-2.5-Pro and Claude-Sonnet-4 implement long CoT reasoning (named Thinking Mode) by default.
+For fair comparison, we disabled Thinking Mode in Claude; Gemini-2.5-Pro does not support full disabling Thinking Mode, we set the thinking token budget to the lowest possible value 128.
+and five recent open-source models that can process multiple images, of sizes ranging from 5B to 12B: Phi-4-Multimodal-Instruct (5.6B, Microsoft 2025), Qwen2.5-VL-7B-Instruct (Bai et al.
+2025), InternVL-3-8B (Zhu et al.
+2025), Llama-3.2-11B-Vision-Instruct (Meta Llama Team 2024), and Pixtral-12B (Mistral AI 2024).
+We use the default inference parameters given by each model in generation.
+CLIP @1 @5 @10 @20 @30
+Recall 2.81 10.26 16.70 25.54 33.16
+NDCG 24.33 25.97 30.98 39.21 46.25
+Hit 24.33 53.74 67.65 77.54 82.62
+Hit Count 0.24 1.01 1.84 3.18 4.38
+BGE @1 @5 @10 @20 @30
+Recall 3.63 12.43 18.81 28.64 35.97
+NDCG 26.74 29.15 33.4 41.83 48.84
+Hit 26.74 58.56 68.98 79.14 83.96
+Hit Count 0.27 1.17 2.04 3.43 4.60
+VLM2Vec @1 @5 @10 @20 @30
+Recall 2.91 10.01 16.03 26.37 33.84
+NDCG 24.87 26.33 30.09 39.13 45.93
+Hit 24.87 51.87 62.83 78.34 82.35
+Hit Count 0.25 1.10 1.91 3.33 4.49
+Table 2: Retrieval results using the sub-corpus of 200-300 images for each query. Even within the small, species level corpus, the model struggles with our challenging text-to-image retrieval task.
+5.2 Results of Text-to-Image Retrieval
+Model Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
+Baselines
+Zero-shot (no image) 35.16 38.90 39.17 41.71 32.35 53.74 60.43 54.28
+GT clue (1 image) 45.04 41.79 43.69 47.11 47.81 59.81 62.88 56.79
+Non-clue (1 image) 34.76 30.08 29.28 42.74 40.37 14.97 17.11 21.39
+Top-K Retrieval-Augmented Generation
+k=1 39.17 41.98 37.57 42.11 44.25 24.06 32.22 35.03
+3 39.30 44.39 41.18 39.71 46.79 41.18 48.53 45.32
+5 37.70 46.93 39.97 41.71 43.80 47.86 54.95 50.40
+7 41.44 45.99 42.65 41.44 43.98 51.34 55.35 52.14
+10 41.44 48.26 42.11 42.65 43.42 49.73 57.62 53.88
+15 41.04 49.73 41.04 44.39 44.92 50.80 60.03 55.48
+20 41.31 50.53 41.31 - - 52.67 61.50 57.35
+One-in-K Augmented Generation
+k=3 41.92 46.85 41.06 47.11 46.73 48.95 59.57 48.05
+5 40.13 48.04 38.44 46.68 45.19 53.22 60.70 49.86
+7 40.80 47.78 39.76 46.68 44.20 55.36 61.53 51.01
+10 39.08 47.43 38.50 44.83 41.90 56.25 62.70 51.69
+15 40.40 49.57 39.14 43.95 41.12 56.36 64.01 52.67
+20 40.83 48.55 39.09 - - 56.50 63.47 52.14
+Table 3: Main experiment results.
+The coloured cells shows the difference with zero-shot score, pink cells indicate performance under zero-shot baseline, light green cells indicate performance over zero-shot, but lower than GT, and green cells indicate outperforming GT.
+All models benefit from ground-truth clue image as augmentation.
+Overall difficulty.
 Successfully retrieving clue images that contain the required visual knowledge is critical to the overall performance of RAG systems in answering the questions.
-The hit rate highlights the difficulty of text-to-image retrieval in our Visual-RAG benchmark.
-Even within the small sub-corpus of 200 - 300 images, there is still 22% chance that no clue image can be retrieved within the top-30 samples.
-The NDCG score also indicates the challenge of this task, i.e., the retrieved clue images are not assigned top similarity scores comparing to the hard negative images.
-6.2 Retrieval-Augmented Generation
-Model Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
-Baselines
-Zero-shot 27.57 32.04 26.17 30.21 35.80 47.21 42.78 42.43
-Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.19
-Non-clue 27.77 33.27 31.53 34.94 31.63 16.61 6.25 8.92
-Retrieved Top-K
-K=1 33.47 39.24 34.13 37.45 34.59 28.09 19.42 18.53
-3 31.14 38.05 32.93 39.44 36.72 35.26 45.48 28.49
-5 35.66 40.24 35.52 37.25 39.51 40.84 45.35 30.88
-7 33.67 40.44 35.33 37.25 36.65 43.82 46.81 35.86
-10 33.47 39.04 37.18 36.85 37.18 44.02 48.40 38.84
-15 33.86 39.64 34.39 37.25 35.99 44.42 51.20 45.62
-20 32.27 40.84 36.59 - - 47.61 53.19 42.23
-Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.79
-One-in-K
-K=3 33.03 40.72 36.72 37.21 37.00 43.90 45.64 32.87
-5 33.97 39.68 35.22 39.70 35.92 45.90 46.35 35.98
-7 33.35 39.36 34.80 35.66 36.53 45.31 45.46 36.60
-10 32.27 39.16 35.61 36.57 34.51 46.14 46.90 37.63
-15 32.23 39.40 34.99 37.32 34.59 46.54 46.58 37.86
-20 32.39 37.73 34.08 - - 45.62 49.57 35.43
-Oracle 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.79
-Table 3: Main experiment results. The coloured cells shows the difference with oracle score, best performing $k$ values for each model is highlighted with boldface.
+As shown in Table 2, across the three retrievers, performance is tightly clustered, with BGE slightly ahead.
+Even when restricting the search space to 200–300 images per species, retrieval remains challenging: across all retrievers, there is $\sim$ 17-18% chance that no clue image appears within the top‑30.
+These results indicate that (i) clue images are genuinely rare and (ii) clue images, if retrieved, are not ranked top by similarity scores.
+The sub-optimal performance of retrievers poses a further challenge for MLLMs, requiring models to tackle with the task under noisy image context.
+Takeaway.
+Substantial performance gap remains to be closed for cross-modal retriever.
+Performance remains far from saturated even at $k=30$, downstream gains will depend not only on the MLLM’s visual reasoning but also on the retriever exposing more clues, and earlier in the ranking.
+5.3 Results of Visual Retrieval-Augmented Generation
 Does image as augmentation benefit MLLMs?
-To assess whether image can be leveraged as evidence in the RAG system, we conduct baseline experiments and compare performance under three conditions: 1) Zero-shot: directly prompting MLLMs with questions without retrieval; 2) Oracle: augmenting the question with a gold clue image; 3) Non-clue: augmenting each question with a non-clue image within the same species, which does not contain visual knowledge relevant to the question [^4] For 2) and 3), the results reported are the average of 5 runs with different clue / non-clue images selected as augmentation., as shown in the first section of Table 3.
-All five open-sourced models exhibit performance gains of up to 15 points when using oracle image as an augmentation.
-This indicates the necessity of including relevant visual knowledge to answer the questions.
-We conclude that for open-sourced models, a single relevant image can serve as good evidence to enhance RAG performance.
-Whereas interestingly, oracle performance for proprietary models drops below the zero-shot setting.
-We hypothesize that the single-image clue provides insufficient evidence to establish the correct visual relationship, causing the model to respond conservatively with “I cannot decide”.
-These conservative responses count as failing to produce an answer, thus reducing the overall scores.
-Specifically, such conservative response occurs in 15.8% of samples for GPT-4o, 7.4% for Gemini, and 16.4% for Claude.
-Meanwhile, results for the Non-clue setting indicate that the models effectively identify these non-clue images as not containing visual knowledge.
-Specifically, all proprietary models frequently respond with “I don’t know” with non-clue images, leading to a substantial drop in scores comparing to the zero-shot and oracle setting.
-Can MLLMs distinguish the clue from retrieved hard negatives?
+To assess whether image can be leveraged as evidence in the RAG system, we conduct baseline experiments and compare performance under three conditions: 1) Zero-shot: directly prompting MLLMs with text questions without image retrieval; 2) GT clue: augmenting the question with one ground-truth clue image; [^4] Results reported are the average of 5 runs with different random clue image selected when possible.
+3) Non-clue: augmenting each question with a non-clue image within the same species, which does not contain visual knowledge relevant to the question, as shown in the first section of Table 3.
+GT clue setting establishes an upper bound on how images can contribute in RAG.
+Relative to zero-shot baseline, supplying a single ground-truth clue image raises performance for all models tested.
+Open-source models receive up to 15 accuracy scores gain, confirming that augmenting relevant visual evidence can substantially compensate for gaps in their pre-trained prior knowledge.
+The three proprietary models, while having strong zero-shot baseline performance, also improved with augmenting clue image for 2-6 points in accuracy.
+Non-clue image, conversely, degrades performance of most models, underscoring that the benefit comes from evidence encoded in the clue itself, not from the mere presence of an image.
+Takeaway: A single, representative clue image reliably improves performance, confirming visual evidence in our benchmark is informative and influential for RAG.
+Does realistic visual RAG benefit MLLMs?
 We have shown that a relevant clue image can serve as good evidence in augmentation.
@@
 For RAG system in real implementation where a number of possibly irrelevant images are retrieved, does retrieval still help the final QA performance?
-To explore this realistic RAG scenario, we provide the top-$k$ retrieved images, derived from Section 6.1, as augmentation to the MLLMs [^5] For pixtral and Llama3.2-V, $k=20$ was not evaluated due to hardware limitation..
+To explore this realistic RAG scenario, we provide the top-$k$ retrieved images, [^5] For Pixtral and Llama3.2-V, $k=20$ was not evaluated due to hardware limitation.
+derived from Section 5.2, as augmentation to the MLLMs.
 The retrieved images are sorted in descending order based on their similarity scores.
-As shown in the second section of Table 3, the open-source models generally lag behind their single-oracle scores, indicating limited multi-image processing capabilities.
-Most open-source models struggle to isolate the clue image and effectively leverage its visual information when it is mingled with hard negatives.
-In contrast, all three proprietary models exhibit a generally monotonic increase in performance as $k$ increases, except Claude which decreases at $k=20$.
-Notably, the performance at $k=5$ begins to surpass oracle baselines, continuing to improve along with higher $k$ values.
-We attribute this improvement to the comparative evaluation enabled by providing both clue and non-clue images.
-Compared to a single clue image, the top-$k$ RAG setting allows the models to contrast relevant and irrelevant visual features, leading to greater confidence in extracting knowledge from the clue images, whereas a single clue image may be insufficient to fully establish its relationship to the question.
-However, the proprietary models fail to exceed their zero-shot scores until fairly large $k$ values, suggesting that interpreting complex visual cues remains challenging, and they may even struggle with reconciling wrongly interpreted visual knowledge with their prior knowledge gained through pre-training.
-As indicated by the retrieval results, even at $k=20$, there is 27% chance that no clue image is included in the retrieved set.
-This likely explains the low QA performance at smaller $k$ values, where failing to retrieve the clue image entirely negates any benefit of visual augmentation.
-To investigate further, we conduct an additional experiment, constructing image sets in which exactly one clue image and $k-1$ non-clue images are provided [^6] Both clue and non-clue images were randomly selected across 5 different combinations, results reported are the average of 5 runs for each $k$..
-The clue image is positioned at the beginning of the $k$ images to simulate the perfect retrieval scenario.
-Demonstrated in the last section of Table 3, the open-sourced models now exhibit decreasing performance as $k$ increases, further strengthen the conclusion of their limited capacity for handling multiple images simultaneously.
-By contrast, all three proprietary models surpass their single-image oracle baseline from $k=3$ onward and maintain near-constant performance at higher $k$ values.
-This suggests that proprietary models are capable of isolating the clue image from irrelevant images, regardless of the number of input images [^7] Additional experiments on randomly positioning the clue image within the set is also carried out.
-The proprietary models keeps the flat performance curve against $k$, indicating there is no positional bias..
-Model Zeroshot Score Augmenting Text
-Phi3.5-V 27.57 26.02
-Qwen2VL 32.04 25.61
-InternVL2.5 26.17 19.70
-Pixtral 30.21 19.74
-Llama3.2-V 35.80 29.73
-Table 4: Results when augmenting retrieved Wikipedia text as evidence.
-6.3 Ablation
-Can the answers for Visual-RAG be found in Wikipedia?
-We construct a small textual knowledge base from Wikipedia articles for the 10k species in iNat21 dataset.
-This minimum knowledge base comprising approximately 97,000 chunks, each consisting of $\sim$150 tokens.
-The detailed collection process is described in Appendix B.
-For text retrieval, we use a pre-trained dense text retriever, DRAGON Plus (Lin et al., 2023).
-Given that the ground-truth labels for relevant text chunks are not available, we only assess whether the retrieved text chunks correspond to the correct Wikipedia page, and the retriever achieves HitRate@10 of 97%.
-Next, we present the QA performance across open-sourced models augmented with top-10 retrieved textual knowledge chunks.
-As shown in Table 4, augmenting the question with textual knowledge from Wikipedia does not improve answer quality for most models.
-Instead, in numerous cases, models fail to locate relevant information within the augmented text and output “I don’t know”.
-This behaviour in turn leads to scores lower than the zero-shot setting.
-These experiments provide strong evidence that the knowledge required to answer the questions related to visual features in our benchmark is unlikely to be contained in textual knowledge bases.
-Is More Clues Always Better?
-We observed that for proprietary models, top-$k$ scores at higher $k$ values excel their 1-in-$k$ score.
-Similarly, the open-source models also achieve their best performance at larger $k$ value in top-$k$ setting than in 1-in-$k$.
-A likely explanation is that the expected number of clue images retrieved in top-$k$ is substantially higher than when only a single clue image is guaranteed.
-We in turn ablate with effect of multiple clue images.
-Fixing $k=10$, we experiment with varying numbers of clue images with $n\in\{1,2,3,4,5\}$ ($n$-in-10).
-We select queries that contain at least 5 clues to enable this experiment.
-As shown in Figure 3, all models exhibit increasing performance as $n$ grows, aligning with the intuition that additional clue images provide more robust visual evidence.
-Figure 3: Impact from number of clue images
-Hard-negatives can be helpful; how about easy-negatives?
-As an ablation to the previous experiment on whether multiple images benefit the overall QA performance, we now mix one clue image with $k-1$ highly irrelevant images from subsets that are not the query species.
-The results indicate that both open-source and proprietary models benefit more from hard negatives coming from the same species compared with easy negatives coming from different species.
-This further strengthen our conclusion that hard negatives are helpful in contrasting the queried features to inform the correct answer.
-Detailed results can be found in Table 9 in Appendix.
+Intuitively, more retrieved images would lead to better performance, as probability of retrieving clue image increase with $k$.
+As shown in the second section of Table 3, while Qwen and the three proprietary models follow the expected trend, other open-source models’ performance saturates and plateaus at medium $k$ (5 - 10), including more retrieved images offers little benefit and sometimes slightly hinders the performance.
+Two intertwined factors plausibly explain the divergence: the open-source models have limited multiple-image processing capacity, as $k$ increase beyond 15, they struggle with effectively extracting related visual information; in the meantime, while larger $k$ brings higher probability of retrieving clue images, the number of negative non-clue images occupy larger portion of the image set.
+This further dilutes the clue signal and magnifies the impact of their restricted multi-image understanding ability.
+Takeaway: Under realistic RAG setting, most of the evaluated models fall noticeably short of the GT clue upper bound, revealing current MLLMs’ limitation in efficiently utilizing retrieved image context.
+However, Qwen stands as a powerful exception, possessing robust multimodal RAG ability.
+Does a guaranteed clue outweigh the noise of additional non-clue images?
+We notice that top-$k$ RAG underperform the GT clue setting for most models.
+Such results conflate two error sources: (i) the retriever may fail to return any clue image, or (ii) clue present, but models fail to extract information alongside visually similar non-clue distractors.
+To investigate further, we conduct an additional experiment named 1-in-$k$, constructing image sets in which exactly one clue image and k1k-1 non-clue images are provided,6665 runs for different random images combinations at each $k$.
+isolating the impact from imperfect retrieval.
+The clue image is positioned at the beginning of the $k$ images, emulating perfect retrieval scenario.
+Demonstrated in the last section of Table 3, most open-source models except Qwen now lose accuracy as $k$ increases.
+Proprietary models, in stark contrast, keep improving monotonically with $k$, while GPT-4o and Gemini surpass their top-$k$ RAG performance.
+Hence, the plateau observed in earlier RAG experiments for open-source models is not merely a retrieval problem: even when the clue is guaranteed, models are still prone to more distractors.
+Conversely, the unsatisfactory performance of proprietary models at lower $k$ under top-$k$ RAG can be confirmed to be attributable from missing clue image retrieved.
+Takeaway: With a guaranteed clue image, proprietary models remain robust to additional non-clue distractors, underscoring their ability to extract visual evidence; whereas open-source models are markedly distractor-sensitive.
+Figure 3: gCUE metrics by models and $k$.
+How efficiently do MLLMs utilize clue and how robust are they against distractors?
+With the 1-in-$k$ results, we propose a quantitatively metric named generalized clue utilization efficiency (gCUE).
+Specifically, we integrate zero-shot and non-clue baselines via $B^{(\lambda)}=\lambda A_{Z}+(1-\lambda)A_{NC}$, where $\lambda\in[0,1]$, $A_{Z}$ denotes zero-shot accuracy, $A_{NC}$ denotes non-clue accuracy, $A_{GT}$ denotes GT clue accuracy and $A_{k}$ denotes 1-in-$k$ accuracy at $k$.
+The non-clue baseline $A_{NC}$ is introduced as an approximation of the impact on performance from non-clue images.
+We define:
+$\mathrm{gCUE}(k;\lambda)=\frac{A_{k}-B^{(\lambda)}}{A_{GT}-B^{(\lambda)}}.$
+$\mathrm{gCUE}$ therefore quantifies the remaining benefit of the clue image over baseline among k1k-1 distractors, compared with sole clue image scenario.
+It describes how effectively the model utilizes the visual clue and how it is prone to distractors.
+$\mathrm{gCUE}=1$ implies model attaining full GT clue benefit, $\mathrm{gCUE}<1$ indicates under-utilization of the clue and vulnerability to distractors, and $\mathrm{gCUE}>1$ suggests model gaining synergy with non-clues.
+By default we set $\lambda=0.5$.
+As shown in Figure 3, Qwen demonstrates strong synergy and distractor tolerance across $k$.
+Other open-source models generally display a trend of decline.
+Their ability to exploit the single clue erodes, suggesting limited capacity for multi-image aggregation and greater vulnerability to a rising distractor load.
+Whereas proprietary models start with low $\mathrm{gCUE}$ at small $k$, implying that a handful of non-clue images can initially overwhelm the clue despite being positioned at front.
+Yet they surprisingly recover as more non-clue context being provided.
+One plausible explanation is that the proprietary models and Qwen implicitly contrast relevant and irrelevant visual features among the noisy image set.
+As more non-clues are added, the distinctive features of the clue become easier to isolate, eventually lead to confidence in extracting knowledge from the clue image.
+A detailed causal analysis of this effect is beyond the scope of the present paper and we leave it to future work.
+Takeaway: $\mathrm{gCUE}$ metric provides an interpretable metric apart from raw accuracy.
+Open-source models except Qwen continue to exhibit inverted dynamics comparing to proprietary models on clue utilization efficiency.
 6 Conclusion
-In this paper, we present Visual-RAG, a challenging VQA benchmark that requires visual knowledge as evidence to answer questions.
-Different from previous works that include images as part of queries and let models retrieve textual knowledge, our Visual-RAG is formed with text-only queries and enforces text-to-image retrieval for challenging visual knowledge.
-We evaluate 8 popular MLLMs under various multimodal RAG settings.
-The results indicate that images can serve as knowledge sources for answering visual knowledge intensive questions, but the models struggle to effectively extract visual knowledge from a single clue image.
-Surprisingly, the current models are able to reason across clue and non-clue images for better understanding of the questions.
-As one of the pioneer works in visual-knowledge centric multimodal RAG, our benchmark can help promote the research and development of MLLMs in extracting fine-grained visual knowledge and enhance cross-modal retrieval in a more realistic setting.
-Limitations
-While our Visual-RAG addresses the lack of visual knowledge intensive QA benchmarks, it has several design constraints.
-First, our query generation and manual rewriting process focus on single-hop questions only.
-Although a small subset of queries might involve simple comparisons within the same image, such as checking the colour of a bird’s upper beak vs.
-lower beak, there are no truly multi-hop queries requiring reasoning across multiple images.
-Designing such multi-image or even cross-species queries would demand significantly more expert knowledge, and we leave this as an avenue for future work.
-Secondly, Visual-RAG focuses exclusively on organisms, which raises questions about the broader generalizability of our findings.
-We chose the domain of organisms because they allow for factual queries regarding visual features that remain consistent across multiple individuals of a species, yet are present in only a minority of images.
-Other commonly used domains in knowledge-intensive VQA such as buildings, artworks, news event, etc., do not always meet these criteria, as answers may not hold universally for all instances or there exists only one instance, and the required visual features are often too prevalent or too trivial to foster a genuinely challenging retrieval scenario.
-Lastly, the overall scale of Visual-RAG is limited compared to larger benchmarks like InfoSeek and Encyclopedic-VQA.
-Their heuristic-driven methods for automatically generating queries and answers do not readily apply to our specialized setting, which hinges on expert-informed checks to ensure the presence (or absence) of specific visual features.
-Despite these constraints, Visual-RAG represents an important step toward truly visual knowledge intensive QA, and we hope it will inspire further research into more complex, multi-hop, and cross-domain tasks.
-Ethical Considerations
+In this paper, we present Visual-RAG, a visual knowledge-intensive VQA benchmark that requires models extracting evidence encoded in images to answer questions.
+Different from previous works that include images as part of queries and let models retrieve textual knowledge, our Visual-RAG is formed with text-only queries and enforces text-to-image retrieval for challenging visual knowledge grounded question answering.
+We evaluate eight popular MLLMs under various multimodal RAG settings.
+The results indicate that images can serve as knowledge sources for answering visual knowledge-intensive questions, but the models struggle to effectively extract visual knowledge from noisy retrieved contexts with hard negatives introduced.
+We further quantify model behavior with a unified efficiency metric $\mathrm{gCUE}$, measuring models’ robustness at different numbers of images.
+As one of the pioneer works in visual-knowledge centric multimodal RAG, our benchmark can help promote the research and development of MLLMs in extracting fine-grained visual knowledge and enhance cross-modal retrieval in a more realistic RAG setting.
+Ethical Statement
 Licensing
@@
 For other potential risks regarding images in iNaturalist 2021 dataset, please refer to the original paper and iNaturalist website (www.inaturalist.org).
-Appendix A Relation with previous works
-Knowledge intensive VQA benchmarks focusing textual knowledge, including but not limited to: ViQuAE (Lerner et al., 2022), InfoSeek (Chen et al., 2023) and Encyclopedic-VQA (Mensink et al., 2023), are discussed in related works.
+Appendix A Extended Comparison with Related Works
+Knowledge intensive VQA benchmarks focusing textual knowledge, including but not limited to: ViQuAE (Lerner et al.
+2022), InfoSeek (Chen et al.
+2023) and Encyclopedic-VQA (Mensink et al.
+2023), are discussed in related works.
 These benchmarks contain questions paired with images of certain entities, and knowledge for such entities can be found in an external textual knowledge base (e.g., Wikipedia).
@@
 Typically, given a question and its associated image, a model is expected to: 1) identify the entity depicted in the image, 2) retrieve relevant documents or passages from a knowledge base that pertain to both the question and the entity, and 3) generate an answer based on the question, image, and retrieved documents.
-Step 1) and 2) is usually carried out by leveraging vision-language pretrained encoders, such as CLIP (Radford et al., 2021), to facilitate cross-modal (image-to-text) retrieval.
+Step 1) and 2) is usually carried out by leveraging vision-language pretrained encoders, such as CLIP (Radford et al.
+2021), to facilitate cross-modal (image-to-text) retrieval.
 While the above mentioned benchmarks are all claimed to be knowledge-intensive VQA benchmarks, it is noteworthy that the knowledge to be retrieved is textual knowledge.
@@
 While cross-modal retrieval was commonly used, the knowledge retrieval process can actually be achieved in uni-modal manner.
-A recent work (Yan and Xie, 2024) achieved “SoTA” performance on InfoSeek and E-VQA employing image-to-image retrieval within a knowledge corpus comprising Wikipedia pages and their associated images.
+Yan and Xie (2024) achieved “SoTA” performance on InfoSeek and E-VQA employing image-to-image retrieval within a knowledge corpus comprising Wikipedia pages and their associated images.
 Entity recognition is performed by identifying the most visually similar Wikipedia image to the query image, which automatically enables retrieval of the related knowledge document (i.e., the Wikipedia page containing that image).
 Answers are then generated using a text-only LLM, augmented with the relevant paragraph from the Wikipedia page, which is extracted through fine-grained reranking.
-WebQA dataset (Chang et al., 2022) included questions that require retrieving images and utilizing visual knowledge to generate answers.
+WebQA dataset (Chang et al.
+2022) included questions that require retrieving images and utilizing visual knowledge to generate answers.
 However, its retrieval process is heavily reliant on image captions, as the images themselves typically lack indicators identifying them as clue images for answering a question.
 For instance, considering the question: “Are the land dinosaurs guarded by rail in both the Display Museum of Natural History in University of Michigan and the Museo Jurassic de Asturias?” Without captions, the model would be unable to discern which dinosaur fossil image corresponds to which museum.
-Table 5 shows the retrieval scores in our implementation of retrieval experiments on WebQA benchmark for queries having images as knowledge source.
+Table 4 shows the retrieval scores in our implementation of retrieval experiments on WebQA benchmark for queries having images as knowledge source, evaluated with clip-vit-large-patch14-336.
 While the queries are text-only, aligning with our setting, the text-to-image retrieval scores significantly lag behind text-to-caption retrieval, especially for lower $k$.
 Recall @1 @5 @10 @20 @50
-Text-to-Image 15.17 37.08 47.63 58.43 72.32
-Text-to-Caption 43.25 68.74 76.86 83.79 89.92
+T2I 15.17 37.08 47.63 58.43 72.32
+T2T 43.25 68.74 76.86 83.79 89.92
 Table 4: Retrieval scores for WebQA dataset.
-MRAG-Bench (Hu et al., 2024) shares a similar objective to ours by proposing a benchmark that requires retrieving visual knowledge to solve questions.
-However, despite this similarity, the focus remains on entity recognition.
-Most tasks are structured such that an image displaying an incomplete or partially visible entity is provided, and the retrieval system is tasked with identifying images that depicts the entire entity.
-This approach primarily enhances the MLLM’s ability to recognize entities rather than addressing broader knowledge-intensive tasks.
-Furthermore, non-clue images in MRAG-Bench are from different entities, which are visually distinct, in contrast to the numerous hard negatives in our Visual-RAG .
-Concurrent to our work, the iNaturalist team released 2024 version of their natural image dataset, iNat24 (Vendrow et al., 2024), coming with a text-to-image retrieval benchmark named INQUIRE, containing 250 queries.
-Though, the queries in INQUIRE are designated for image retrieval only, and are not QA style questions that have explicit answers.
-As iNat24 and iNat21 share the same taxonomy classes of organisms, we do consider adding the new images in iNat24 to our benchmark.
-We also note a related line of research (Tanaka et al., 2023; Van Landeghem et al., 2023) on extracting knowledge from document images (e.g., PDF files, Powerpoint slides), conventionally referred to as OCR-based VQA.
+T2I: Query-to-Image retrieval.
+T2T: Query-to-Caption retrieval.
+T2T retrieval significantly outperform T2I, indicating crucial identifier for images being the clue reside in caption.
+In addition to MRAG-Bench introduced in the main paper, another recent work M2RAG (Liu et al.
+2025) gathers data samples from existing benchmarks, namely WebQA introduced in previous paragraph and Factify (Mishra et al.
+2022), a multimodal fact verification dataset.
+The multi-modal QA subtask introduced partially overlap with ours, while the queries are repackaged instances drawn from the original WebQA dataset.
+The iNaturalist team released 2024 version of their natural image dataset, iNat24 (Vendrow et al.
+2024), coming with a text-to-image retrieval benchmark named INQUIRE, containing 250 queries.
+Though, the queries in INQUIRE are designated for image retrieval only instead of QA style questions that have explicit answers.
+As iNat24 and iNat21 share the same taxonomy classes of organisms and contain newly uploaded images not included in iNat21, we do consider adding the new images in iNat24 to our benchmark.
+We conduct preliminary experiment using iNat24 images as retrieval corpus, evaluated with Qwen2.5VL, as shown in Table 5.
+The overall performance remains at comparable level.
+Top-$k$ RAG iNat21 iNat24
+k=1 41.98 42.72
+k=3 44.39 42.18
+k=5 46.93 45.96
+k=7 45.99 46.23
+k=10 48.26 45.28
+k=15 49.73 48.47
+k=20 50.53 48.79
+Table 5: Top-$k$ RAG performance using iNat24 images as knowledge corpus, evaluated with Qwen2.5VL.
+We also note a related line of research (Tanaka et al.
+2023; Van Landeghem et al.
+2023) on extracting knowledge from document images (e.g., PDF files, Powerpoint slides), conventionally referred to as OCR-based VQA.
 Although these tasks also require models to retrieve relevant page and interpret images to answer textual queries, the underlying knowledge is primarily textual—obtained via OCR.
-This focus diverges considerably from our goal of visual feature extraction, where visual cues are inherently sparser in the wild.
-Appendix B Construction of Minimum Textual Knowledge Base
+This focus diverges considerably from our goal of visual feature extraction, where visual cues are inherently more sparse in the wild.
+Appendix B Additional Ablation Studies
+B.1 Can the answers for Visual-RAG be found in Wikipedia?
+We construct a small textual knowledge base from Wikipedia articles for the 10k species in iNat21 dataset.
+This minimum knowledge base comprising approximately 97,000 chunks, each consisting of $\sim$150 tokens.
 Among the 10000 taxonomy species in iNat21 dataset, we matched 9614 Wikipedia pages.
@@
 Each chunk is bounded within a single section of the original article to preserve contextual integrity.
-Sections such as “Reference”, “See also”, “External links”, etc., which lack informative content, were excluded.
-The chunks are cut upon period mark, so that they won’t end or start with incomplete sentence.
-Two adjacent sections, having total length less than 150 tokens will be merged, with the section titles kept.
+Sections such as “Reference”, “See also”, “External links”, etc., which lack informative content were excluded.
+The chunks are cut upon period mark, so that they do not end or start with incomplete sentence.
+Two or more adjacent sections having total length less than 150 tokens are merged, with the section titles kept.
+For text retrieval, we use a pre-trained dense text retriever, DRAGON Plus (Lin et al.
+2023).
+Given that the ground-truth labels for relevant text chunks are not available, we only assess whether the retrieved text chunks correspond to the correct Wikipedia page, and the retriever achieves Hit@10 of 97%.
+Next, we present the QA performance across open-sourced models augmented with top-10 retrieved textual knowledge chunks.
+As shown in Table 6, augmenting with textual knowledge from Wikipedia result in performance drop for both models tested.
+Instead, in numerous cases, models fail to locate relevant information within the augmented text and output “I don’t know”.
+This behaviour in turn leads to scores lower than the zero-shot setting.
+These experiments provide strong evidence that the knowledge required to answer the questions related to visual features in our benchmark is unlikely to be textually documented in Wikipedia.
+Qwen2.5VL GPT-4o
+Zero-shot 40.23 56.58
+RAG-Best 51.88(@20) 56.58(@7)
+Text RAG 25.53 31.68
+1GT 41.73 59.89
+2GT 48.12 53.38
+3GT 54.51 54.89
+4GT 53.20 56.77
+5GT 52.44 60.15
+1-in-10 45.90 56.94
+2-in-10 47.20 58.08
+3-in-10 51.13 61.09
+4-in-10 52.19 59.21
+5-in-10 52.07 61.47
+Table 6: Ablation experiment: results when augmenting textual knowledge and impact from number of clue images.
+1GT - 5GT: input 1-5 ground truth clue images only (mmGT setting); 1-in-10: 1 clue with 9 non-clues, 2 clues with 8 non-clues, etc.
+B.2 Does including more clue images consistently lead to better performance?
+We observed that in most models, top-$k$ RAG scores at higher $k$ values excel their 1-in-$k$ score.
+A likely explanation is that the expected number of clue images retrieved in top-$k$ is substantially higher than when only a single clue image is guaranteed.
+We in turn ablate with effect of multiple clue images.
+We compare two regimes on the subset of queries with sufficient clues: mmGT ($m$ ground-truth clue images) and $m$-in-10 ($m$ clues mixed with 10m10{-}m non-clues from the same species), reported in Table 6.
+Aligned with main experiments, mixing non-clues is beneficial at low $m$: non-clue context appears to regularize both models’ multi-image aggregation ability, improving robustness when only a few clues are present.
+For Qwen, the mmGT results forms a bell-shaped curve, where 3 clue images gives the best performance, while $m$-in-10 gives overall increasing curve.
+This suggests Qwen benefits most from a small set of clean, diverse clue views, and is sensitive to additional non-clue context.
+GPT-4o, conversely, gives a U-shaped curve for mmGT.
+Interestingly, small number of clue images corrupts the model’s correct answer when only one clue is present, and mixing them with non-clues brings the performance close to optimal.
+Same Species Different Species
+k=3 46.85 46.15
+k=5 48.04 47.41
+k=7 47.78 46.97
+k=10 47.43 47.55
+k=15 49.57 47.72
+k=20 48.55 47.76
+Table 7: Impact from using highly irrelevant distractor in 1-in-k setting, evaluated with Qwen2.5VL.
+B.3 Hard negatives can be helpful; how about easy negatives?
+We observed that for Qwen2.5VL and some proprietary models, the performance of mixing non-clues with a clue image can sometimes surpass the one clue image upper-bound.
+As an ablation to the experiment findings, we now mix one clue image with k1k-1 highly irrelevant images from subsets that are not the query species.
+The results in Table 7 indicate that benefit from hard negatives overall surpass that from easy negatives, especially at higher $k$, partially confirming our hypothesis in last experiment of Section 5.3.
+As an example, for question “What color are the hindwings of Papaipema inquaesita?”, hard negatives displays the moth not showing the hindwings, while easy negatives are images of plants, mammals, etc.
+Without the normal image showing only forewings, model could be confused even when presented the hindwings.
+B.4 Does visual clue reduce “I don’t know” responses?
+During qualitative inspection we observed that, at small $k$, proprietary models often answer with phrases such as “I cannot decide” or “Not enough information.” We conventionally name such outputs as “IDK” responses.
+A natural hypothesis is that adding clue images should increase model confidence and therefore lower the IDK rate.
+Table 8 confirms this for the proprietary models and for Qwen: as $k$ grows, their IDK rate falls, mirroring the rise in accuracy.
+By contrast, Phi, Pixtral, and Llama produce few IDK responses even when their accuracy fluctuates, suggesting that these models tend to commit to an answer rather than abstain.
+This tendency also explains the seemingly counter-intuitive result that Llama’s accuracy in the Non-clue setting exceeds its zero-shot baseline.
+Although the added non-clue image lacks the required visual evidence, it could provide the model with organism-specific context, prompting it to draw on parametric knowledge to answer.
+Model Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
+Baselines
+Zero-shot (no image) 0.27 0.00 1.60 0.53 11.76 0.00 0.53 3.74
+GT clue (1 image) 0.59 9.89 6.52 0.48 2.84 4.60 2.41 5.83
+Non-clue (1 image) 2.14 27.27 31.28 1.87 4.81 76.74 71.12 67.11
+Top-K Retrieval-Augmented Generation
+k=1 0.80 10.43 13.10 0.80 2.67 58.02 47.86 41.18
+3 0.80 8.29 8.29 0.00 1.07 27.54 24.60 20.86
+5 0.53 3.74 8.82 0.53 1.34 18.45 16.58 12.57
+7 0.80 3.74 7.22 1.07 0.53 15.24 14.17 11.23
+10 0.27 2.94 6.42 0.27 1.60 15.24 11.50 8.82
+15 2.14 2.94 5.08 0.27 1.34 13.64 9.36 7.22
+20 1.87 2.14 3.21 - - 10.70 6.15 8.02
+1-in-K Augmented Generation
+k=3 1.34 6.42 10.64 0.32 1.02 21.76 8.18 16.36
+5 1.02 3.21 9.25 0.27 0.85 15.08 5.77 13.10
+7 1.07 2.94 10.64 0.37 1.34 12.41 4.49 10.43
+10 1.18 2.40 10.32 0.22 1.17 11.39 3.96 9.46
+15 2.41 1.39 8.23 0.32 1.39 9.36 2.46 7.33
+20 2.73 1.93 4.97 - - 8.66 1.93 7.65
+Table 8: Percentages of answering “I don’t know” across models and settings.
 Appendix C Annotation details
+For organism species selection in query construction, we aim to ensure that the chosen species are relatively under-represented in textual resources.To achieve that, we filter them based on the length of their Wikipedia summary and description sections.
+Well-known species typically have multiple sections (e.g., habitats, behaviour, etc.), indicating extensive documentation.
+We therefore include only those species whose summary and description account for no more than 50% length of the entire Wikipedia article, as an attempt on minimizing reliance on prior knowledge: widely recognized species may have detailed descriptions on visual features across LLM pre-training corpora, even if these details are not explicitly mentioned in Wikipedia.
 For the query generation, we mentioned that we have gone through a human annotator originated refined filtering and rewriting process.
-The prompt template is shown in Table LABEL:tab:prompt_qgen, and following shows the example prototype queries generated with the sample prompt:
+The prompt template is shown in Table 10, and following shows the example prototype queries generated with the sample prompt:
+Figure 4: Example image of Bromus catharticus flower anther. Owned by original uploader at www.inaturalist.org, shared under $C$-BY-NC license.
   1. 1.
@@
 Query too obscure and likely no clue image can be found.
-Figure 4: Image of Bromus catharticus flower anther. Owned by original uploader at www.inaturalist.org, shared under CC-BY-NC license.
 Human annotator carefully check the candidate prototype queries by browsing all available material online with the scientific name as search keyword, as well as briefly scan through the image corpus by viewing the image thumbnails.
-After query generation and rewriting, and coarse-level filtering described in section 4, we dispatch the validated queries to the next round of human annotation of image labels.
+After query generation and rewriting, and the coarse-level filtering described in Section 4.2, we dispatch the validated queries to the next round of human annotation of image labels.
 Figure 5 shows the annotation interface for image labelling.
@@
 In the meanwhile, annotators provide query answers upon finding a clue image, as shown in Figure 6.
-Annotators are allowed to skip for a query when finding more than 9 clues after the first 27 images, indicating this query likly fail the clue image threshold.
+Annotators are allowed to skip for a query when finding more than 9 clues after the first 27 images, indicating this query likely fail the clue image threshold.
 The order of images are randomly shuffled.
-Figure 5: Annotation interface for image labelling
-Figure 6: Annotation interface for image label and query answer
-Appendix D Prompt Template
-This section list the prompt templates used through out our experiments.
+Figure 5: Annotation interface for image labelling.
+Figure 6: Annotation interface for image label and query answer.
+[…]
+True 86 1 3
+Partial 6 14 2
+False 6 17 65
+Table 9: Confusion matrix of human evaluation v.s. LLM-as-Judge. We consider “partially correct” as wildcard match, and the macro-$F_{1}$ score is 0.92.
+Appendix D Prompt Templates
+Table 10, 11, 12 list the prompt templates used through out our experiments.
 Prompt Template for Query Generation
@@
 Please avoid question regarding visual features explicitly mentioned in the Wikipedia paragraphs.
-(e.g., paragraph: “The seeds are colored in red” -> Don’t ask “What color is the seed”; paragraph: “The plant have symbolic seed pods” (but doesn’t mention its color) -> You may ask “What shape/color is the seed pod”)
+(e.g., paragraph: “The seeds are colored in red” ->> Don’t ask “What color is the seed”; paragraph: “The plant have symbolic seed pods” (but doesn’t mention its color) ->> You may ask “What shape/color is the seed pod”)
 Following are format instructions.
@@
 What is the common color of tip of leaf of wall-screw moss (scientific name: Tortula muralis)?
-(Explanation: The moss can be distinguished from similar mosses by its ḧair-pointedl̈eaves.) (Feature: color of leaf tip)
+(Explanation: The moss can be distinguished from similar mosses by its “air-pointed” leaves.) (Feature: color of leaf tip)
 Scientific name: Ptiliogonys cinereus
@@
 (Feature: color of undertail covert viewing from front)
-2. What pattern can be observed on the tail feather of Gray Silky-flycatcher (scientific name: Ptiliogonys cinereus) from dorsal (back) view? (Feature: pattern on tail feather from dorsal view)
-More ICL Examples omitted
+2.
+What pattern can be observed on the tail feather of Gray Silky-flycatcher (scientific name: Ptiliogonys cinereus) from dorsal (back) view?
+(Feature: pattern on tail feather from dorsal view) … More ICL Examples omitted
 End of examples.
@@
 Common Name: Rescuegrass
-Wiki Paragraph:
-Bromus catharticus is a species of brome grass known by the common names rescuegrass, grazing brome, prairie grass, and Schrader’s bromegrass.
+Wiki Paragraph: Bromus catharticus is a species of brome grass known by the common names rescuegrass, grazing brome, prairie grass, and Schrader’s bromegrass.
 The specific epithet catharticus is Latin, meaning cathartic.
@@
 Question:
-Table 6: Prompt Template for Query Generation. Texts in italic and purple are not part of prompt.
+Table 10: Prompt Template for Query Generation. Texts in italic are not part of prompt.
 Prompt Template for Main Question Answering Experiments of Different Settings
@@
 Question:
-Oracle (one clue)
+GT clue
 Please answer the question regarding a visual feature of an organism (animal, plant, etc.).
@@
 Question:
-<image>
+<<image>>
 Non-Clue
@@
 Question:
-<image>
+<<image>>
 Multiple Images
@@
 Question:
-<image> <image> <image>…
+<<image>> <<image>> <<image>>
 Table 11: Prompt Template for Main Question Answering Experiments of Different Settings.
@@
 The reference answer is provided, and the reference is usually short phrases or a single keyword.
-If the student answer is containing the keywords or similar expressions (including similar color), without any additional guessed information, it is full correct.
+If the student answer is containing the keywords or similar expressions (including similar or close color/pattern), without any additional guessed information, it is full correct.
+Similar or close color/pattern includes but not limited to the following cases: pale or light color can appear to be yellowish/greyish under different light condition; dark colors like dark brown, dark grey, dark purple can appear close to each other, and may appear as black as well; stripe pattern can appear as band or ring, dotted pattern can etc.
 If the student answer have missed some important part in the reference answer, please assign partial score.
-Usually, when there are 2 key features and only 1 is being answered, assign 0.5 score; if there are more than 2 key features, adjust partial score by ratio of correctly answered key feature.
 The reference answer can be in the form of a Python list, in this case, any one of the list item is correct.
-If student answer contain irrelevant information not related to question, mark it with “Redundant”, but it does not affect score if related part are correct.
+If student answer contain irrelevant information not related to question, mark it with “Redundant”, but it does not affect score if related parts are correct.
 (e.g.
-Question: what shape is leave of Sanguinaria canadensis, Student Answer: shape is xxx, color is yyy, this is Redundant answer)
-If student answer contain features not listed in reference answer, mark it with “Likely Hallucination” and deduct 0.5 score.
+Question: what shape are leaves of XYZ plant, Student Answer: shape xxx, color yyy, color is Redundant answer)
+If student answer contain features not listed in reference answer, deduct 0.5 score and mark it with “Likely Hallucination”.
 (e.g., Reference Answer: black and white.
-Student Answer: black white, with yellow dots, “yellow dots” is not mentioned in reference)
-Separate the remarks with score using “|”, that is, use the syntax of: “Score: score | Likely Hallucination”, “Score: {score}”, “Score: {score} | Likely Hallucination | Redundant”, etc.
+Student Answer: black white, with yellow dots, “Yellow dots” is not mentioned in reference).
+The reference answer sometimes contains additional information not asked in question, usually enclosed by brackets (), to help verifying hallucinations (e.g.: “Shape is xxx (color is yyy)”).
+Not mentioning additional information in answer is not considered wrong.
+For yes/no question, reference may contain explanations on why giving yes/no, but it is not necessary for student to answer the explanation; however, if student explains and differs with reference, it is considered as hallucination.
+Answering “I don’t know”, “Not enough information” or similar is considered wrong, and please mark it with “No Answer”.
+Format Instructions: Separate the remarks with score using “||”, that is, use the syntax of: “Score: score $|$ Likely Hallucination”, “Score: score”, “Score: score $|$ Likely Hallucination $|$ Redundant”, “Score: 0 $|$ No Answer ”.
 If any explanation on why giving the score is needed, do not start a new line and append after remark with brackets, e.g.
@@
 Question: Is there any specific color marking around the eyes of a semipalmated plover (scientific name: Charadrius semipalmatus)?
-Reference Answer: black eye-round feather, white stripe above eyes, sometimes connected to the white forehead
+Reference Answer: black eye-round feather, white stripe above eyes. (sometimes connected to the white forehead)
 Student Answer: Yes, the bird has a distinctive black line that runs through the eye, which is a key identifying feature.
@@
 Score: 0.5 $|$ Likely Hallucination (Explanation: not white ring, but only a line above the eye)
-Question: …
-More ICL Examples omitted
+Question: What is the typical color of the antennae of Harris’s checkerspot butterfly (scientific name: Chlosyne harrisii)?
+Reference Answer: alternating black and white band, with yellow on the tip
+Student Answer: The antennae of Harris’s checkerspot butterfly are black with orange-tipped clubs.
+Score: 0.5 (Explanation: not mentioning black and white)
+Student Answer: The typical color of the antennae of Harris’s checkerspot butterfly is black with white spots.
+Score: 0.5 $|$ Likely Hallucination (Explanation: not white spot but band. Not mentioning the tip)
+Question: Are the leaves of burro-weed (scientific name: Ambrosia dumosa) usually covered in small hairs?
+Reference Answer: yes
+Student Answer: Yes, the leaves of burro-weed (Ambrosia dumosa) are typically covered in small hairs, giving them a grayish or whitish-green appearance.
+Score: 1 $|$ Redundant
+… More ICL examples omitted…
 Now, Score the following question:
-Table 8: Prompt Template For Automatic Evaluation.
-Note that for evaluation, no images is provided, and it becomes pure-text open-ended QA evaluation.
-Texts in italic and purple are not part of prompt.
+Table 12: Prompt Template For Automatic Evaluation. Note that for evaluation, no images is provided, and it becomes pure-text open-ended QA evaluation. Texts in italic are not part of prompt.
 Appendix E Detailed Results
-This section contains additional experimental results.
-We also include ROUGE-1 recall scores [^9] Implemented with https://pypi.org/project/rouge/, version 1.0.1 for readers who are interested.
-However, because ROUGE relies on strict word matching, the results for open-ended QA can vary significantly due to the different generation styles across models.
-Model Qwen2VL Llama3.2-V GPT-4o
-Same Irrel Same Irrel Same Irrel
-k=3 40.72 36.27 37.00 36.75 43.90 39.79
-5 39.68 37.53 35.92 32.16 45.90 45.03
-7 39.36 35.66 36.53 29.55 45.31 40.36
-10 39.16 33.30 34.51 27.80 46.14 45.29
-15 39.40 35.56 34.59 28.19 46.54 45.64
-20 37.73 32.12 - - 45.62 44.50
-Table 9: Impact from using highly irrelevant distractor in 1-in-k setting. “Same” denotes hard-negatives from same species, “Irrel” denotes irrelevant images from other species.
-Setting Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
-Zeroshot 27.57 32.04 26.17 30.21 35.80 47.21 42.78 42.43
-Oracle 33.67 44.82 38.38 35.96 43.89 38.45 19.6 33.67
-34.26 43.63 37.05 37.25 42.82 39.24 18.49 32.47
-30.88 43.82 39.37 39.44 37.12 37.25 18.22 27.69
-32.67 40.64 39.97 39.04 45.08 36.25 19.55 33.47
-33.07 43.03 38.11 37.25 38.65 37.05 16.94 33.67
-AVG 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.19
-STD 1.285 1.563 1.136 1.433 3.449 1.188 1.096 2.566
-Non-clue 26.69 33.27 31.21 33.27 30.28 14.74 4.46 8.17
-27.49 34.26 34.06 35.86 34.66 17.73 7.41 8.76
-30.28 33.86 32.07 37.05 29.88 17.93 6.71 10.16
-27.09 33.47 29.68 34.66 32.53 16.53 7.07 8.96
-27.29 31.47 30.61 33.86 30.81 16.14 5.59 8.57
-AVG 27.77 33.27 31.53 34.94 31.63 16.61 6.25 8.92
-STD 1.435 1.073 1.663 1.527 1.972 1.296 1.211 0.750
+This section contains detailed experimental results.
+The breakdown results for main experiments can be found in Table 13, 15.
+We also include ROUGE-1 Recall scores [^8] Implemented with https://pypi.org/project/rouge/, version 1.0.1 for readers who are interested in Table 14 and 16.
+However, ROUGE relies on strict word matching, the results for open-ended QA can vary significantly due to the different generation styles across models.
+Setting Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
+Zeroshot 35.16 38.90 39.17 41.71 32.35 53.74 60.43 54.28
+Non-clue 34.76 30.08 29.28 42.74 40.37 14.97 17.11 21.39
+GT clue 45.05 42.51 42.11 46.66 46.26 62.17 62.17 55.35
+44.92 38.50 41.44 47.86 45.99 56.95 61.76 60.56
+45.94 43.58 43.32 47.73 47.46 60.16 63.10 55.88
+46.79 43.72 44.92 45.19 50.40 60.29 64.57 55.88
+42.51 40.64 46.66 48.13 48.93 59.49 62.79 56.28
+AVG 45.04 41.79 43.69 47.11 47.81 59.81 62.88 56.79
+STD 1.604 2.213 2.123 1.212 1.857 1.884 1.081 2.133
 Table 13: Baselines detailed results
-1 in K Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
-k=3 32.27 40.44 37.58 37.45 39.04 41.63 42.94 32.27
-32.27 39.64 36.52 35.26 34.79 44.22 49.09 32.67
-34.06 40.24 37.71 37.85 34.53 44.02 47.43 35.26
-33.86 41.63 36.12 39.64 38.84 43.03 41.99 30.68
-32.67 41.63 35.65 35.86 37.78 46.61 46.75 33.47
-AVG 33.03 40.72 36.72 37.21 37.00 43.90 45.64 32.87
-STD 0.871 0.885 0.903 1.732 2.187 1.828 3.039 1.678
-k=5 33.86 39.84 34.39 40.93 37.78 46.81 47.8 37.25
-33.20 39.64 34.46 38.05 33.20 46.02 46.54 34.86
-34.46 40.24 35.45 44.2 35.06 47.01 46.61 36.06
-33.67 39.64 35.99 36.85 37.12 44.42 43.61 35.26
-34.66 39.04 35.79 38.45 36.45 45.22 47.18 36.45
-AVG 33.97 39.68 35.22 39.70 35.92 45.90 46.35 35.98
-STD 0.594 0.434 0.748 2.923 1.825 1.087 1.613 0.950
-k=7 31.08 38.65 35.45 34.66 37.91 43.03 46.25 37.4
-32.27 39.44 34.13 37.05 35.39 44.02 44.97 35.88
-32.27 38.84 36.32 33.47 34.59 45.29 44.71 36.83
-36.85 39.44 35.06 36.65 35.85 46.22 45.66 35.88
-34.26 40.44 33.06 36.46 38.91 48.01 45.71 37.02
-AVG 33.35 39.36 34.80 35.66 36.53 45.31 45.46 36.60
-STD 2.267 0.699 1.253 1.530 1.809 1.935 0.618 0.690
-k=10 32.27 39.04 35.72 37.00 33.53 45.62 46.73 34.15
-33.07 39.64 37.05 37.10 34.45 44.82 48.79 37.65
-32.87 39.84 35.12 37.70 36.39 46.41 49.66 38.25
-31.47 38.45 35.78 36.23 33.20 47.61 43.54 39.24
-31.67 38.84 34.39 34.82 34.99 46.22 45.76 38.84
-AVG 32.27 39.16 35.61 36.57 34.51 46.14 46.90 37.63
-STD 0.707 0.573 0.980 1.109 1.269 1.031 2.440 2.034
-k=15 31.27 40.44 33.86 42.01 31.94 45.42 47.6 34.15
-33.86 39.24 35.99 36.55 34.92 48.01 49.63 37.79
-30.08 40.24 34.66 36.38 35.52 47.81 42.42 38.74
-32.47 38.84 36.65 36.38 36.58 46.02 47.6 39.5
-33.47 38.25 33.79 35.27 34.00 45.42 45.66 39.12
-AVG 32.23 39.40 34.99 37.32 34.59 46.54 46.58 37.86
-STD 1.565 0.929 1.283 2.672 1.754 1.280 2.717 2.169
-k=20 33.47 38.05 33.93 - - 44.82 51.15 35.67
-29.48 39.44 35.32 - - 44.62 47.99 34.76
-33.47 36.25 31.94 - - 45.22 47.13 36.39
-32.47 37.65 34.00 - - 47.21 48.57 33.47
-33.07 37.25 35.19 - - 46.22 53.01 36.85
-AVG 32.39 37.73 34.08 - - 45.62 49.57 35.43
-STD 1.679 1.168 1.358 - - 1.083 2.438 1.350
+Setting Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
+Zeroshot 21.46 27.06 25.66 25.88 16.63 24.00 26.47 35.38
+Non-clue 13.96 12.33 19.81 23.52 3.96 13.76 16.01 24.14
+GT clue 18.79 27.78 20.64 28.11 15.55 25.25 24.34 32.36
+19.29 25.89 21.16 26.74 15.73 26.51 22.44 33.88
+18.74 27.43 21.36 27.68 16.18 25.64 22.94 33.45
+19.70 27.49 22.71 26.64 16.27 24.26 24.15 32.37
+18.91 26.89 24.19 26.48 14.53 26.67 24.38 33.46
+AVG 19.09 27.10 22.01 27.13 15.65 25.67 23.65 33.10
+STD 0.405 0.747 1.437 0.721 0.696 0.984 0.898 0.697
+Top-K Retrieval-Augmented Generation
+k=1 16.03 13.58 21.71 23.58 4.54 14.48 20.57 28.22
+k=3 16.92 7.95 20.66 22.32 7.37 19.60 24.50 32.78
+k=5 14.99 11.45 18.98 21.92 5.46 22.59 25.67 34.04
+k=7 16.15 12.81 17.62 21.51 6.11 24.19 27.92 34.73
+k=10 16.61 13.88 18.40 22.49 6.73 23.49 28.17 35.18
+k=15 16.55 13.55 16.94 24.60 5.72 23.74 31.12 35.43
+k=20 16.75 13.51 14.93 - - 25.50 30.30 37.95
+Table 14: Baselines and Top-K RAG ROUGE scores for reference. Please note that ROUGE score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.
+1 in K Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
+k=3 41.04 46.93 39.44 48.26 46.26 47.99 59.09 48.26
+43.54 47.46 41.67 46.12 46.75 50.22 60.29 48.80
+40.78 45.05 41.18 48.80 48.53 47.33 58.82 47.46
+41.98 46.26 43.85 45.86 47.73 50.27 60.29 47.33
+42.25 48.53 39.17 46.52 44.39 48.93 59.36 48.40
+AVG 41.92 46.85 41.06 47.11 46.73 48.95 59.57 48.05
+STD 1.097 1.303 1.895 1.329 1.576 1.314 0.684 0.632
+k=5 43.45 48.13 37.97 46.39 47.59 54.41 61.50 50.00
+39.30 48.53 40.37 47.59 45.05 51.28 59.49 49.06
+38.50 47.55 36.32 46.52 43.72 54.28 62.57 49.06
+39.44 49.06 38.24 46.66 45.86 52.67 59.36 50.40
+39.97 46.93 39.30 46.26 43.72 53.48 60.56 50.80
+AVG 40.13 48.04 38.44 46.68 45.19 53.22 60.70 49.86
+STD 1.928 0.831 1.518 0.528 1.624 1.292 1.362 0.787
+k=7 40.37 48.80 40.64 46.39 43.32 56.68 61.93 51.20
+41.04 47.06 36.36 47.59 42.51 56.37 61.90 52.81
+39.57 47.46 40.64 46.52 46.79 54.41 61.36 50.48
+40.64 48.26 41.44 46.66 42.38 51.34 61.76 50.94
+42.38 47.33 39.71 46.26 45.99 58.02 60.70 49.60
+AVG 40.80 47.78 39.76 46.68 44.20 55.36 61.53 51.01
+STD 1.034 0.724 1.996 0.528 2.053 2.593 0.517 1.178
+k=10 40.24 45.19 37.43 44.65 41.44 53.61 62.70 50.27
+41.00 47.06 39.44 46.12 37.17 57.49 60.56 51.07
+37.97 48.66 38.64 44.79 42.38 56.23 63.90 53.88
+39.57 49.06 38.50 44.61 42.91 55.75 63.64 53.21
+36.63 47.19 38.50 43.98 45.59 58.16 62.70 50.00
+AVG 40.80 47.43 38.50 44.83 41.90 56.25 62.70 51.69
+STD 1.034 1.531 0.716 0.786 3.060 1.761 1.314 1.758
+k=15 41.31 51.34 39.30 44.92 43.32 57.62 64.17 52.01
+39.57 49.73 37.83 43.72 39.17 55.88 64.71 51.20
+39.71 51.07 38.90 43.58 38.50 55.08 63.90 54.68
+42.25 47.59 39.17 43.54 42.11 57.49 62.37 54.81
+39.17 48.13 40.51 43.98 42.51 55.75 64.88 50.67
+AVG 39.08 49.57 39.14 43.95 41.12 56.36 64.01 52.67
+STD 1.768 1.688 0.959 0.570 2.146 1.130 0.997 1.950
+k=20 41.58 47.33 39.30 - - 56.95 65.37 52.54
+39.84 49.33 37.03 - - 57.22 64.57 50.13
+41.18 50.94 39.44 - - 54.81 62.03 53.21
+39.71 47.15 42.78 - - 55.48 62.43 51.60
+41.84 47.99 36.90 - - 58.02 62.97 53.21
+AVG 37.25 48.55 39.09 - - 56.50 63.47 52.14
+STD 0.992 1.588 2.389 - - 1.316 1.434 1.302
 Table 15: 1-in-K detailed results
-Setting Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
-Zeroshot 14.66 14.87 17.03 20.44 14.05 11.96 12.58 9.65
-Oracle 15.26 26.75 17.16 25.83 6.72 16.48 16.35 24.59
-15.85 26.63 19.01 25.96 8.1 15.9 14.1 23.26
-15.22 25.86 18.72 29.25 7.19 18.44 14.77 23.63
-15.79 25.08 19.46 28.42 10.39 20.07 14.58 23.83
-15.27 26.44 17.06 25.86 7.88 20.36 11.47 26.57
-AVG 15.48 26.15 18.28 27.06 8.06 18.25 14.25 24.38
-STD 0.313 0.690 1.102 1.644 1.416 2.028 1.771 1.319
-Non-clue 9.49 21.59 8.07 23.25 3.95 9.23 9.04 17.11
-9.27 21.13 9.24 23.94 5.4 7.86 10.03 17.18
-8.89 20.96 10.78 24.4 2.87 10.06 9.21 17.42
-9.85 21.23 8.69 24.63 5 6.62 11.4 15.26
-9.27 20.86 9.05 24.18 4.05 9.9 10.14 16.09
-AVG 9.35 21.15 9.17 24.08 4.25 8.73 9.96 16.61
-STD 0.351 0.283 1.006 0.530 0.990 1.466 0.938 0.912
-Table 12: Baselines rouge scores for reference. Please note that rouge score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.
-1 in K Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
-k=3 10.49 23.80 8.79 24.31 6.92 26.23 16.14 23.27
-10.30 25.47 9.65 24.12 6.64 20.81 18.71 22.65
-10.57 23.89 8.22 25.83 5.55 24.87 19.38 22.69
-9.89 24.72 8.78 25.27 7.97 23.78 17.91 22.74
-10.29 25.80 9.34 25.79 6.76 24.18 19.19 21.31
-AVG 10.31 24.74 8.96 25.06 6.77 23.97 18.27 22.53
-STD 0.263 0.903 0.554 0.809 0.861 1.999 1.317 0.728
-k=5 12.18 25.82 7.77 25.82 8.71 22.44 16.94 22.37
-9.87 23.39 8.22 25.45 4.79 21.04 16.36 23.2
-10.50 23.54 8.65 26.76 5.55 25.28 16.95 23.62
-11.76 22.36 7.72 22.11 7.65 17.79 16.17 23.17
-10.30 24.67 8.62 24.56 5.12 28.04 17.84 24.35
-AVG 10.92 23.72 8.20 24.94 6.36 22.92 16.85 23.34
-STD 0.995 0.967 0.446 1.768 1.721 3.933 0.652 0.722
-k=7 9.28 23.26 8.63 21.65 7.57 22.62 15.67 24.38
-10.72 24.28 8.17 24.61 5.66 25.55 20.08 21.73
-11.45 24.74 8.38 21.51 7.14 20.67 16.35 24
-12.10 24.81 8.22 24.40 7.63 25.97 17.62 22.79
-12.48 24.32 7.67 23.57 7.01 24.4 18.82 23.07
-AVG 11.21 24.28 8.21 23.15 7.00 23.84 17.71 23.19
-STD 1.268 0.619 0.353 1.484 0.796 2.197 1.794 1.046
-k=10 11.85 23.90 7.65 25.71 5.81 23.86 18.92 23.03
-12.15 24.00 8.81 25.50 6.93 24.44 17.5 21.05
-11.83 23.69 8.19 24.96 6.15 23.38 18.64 21.52
-11.79 24.45 8.77 25.28 6.62 24.34 17.57 23.31
-11.06 24.63 8.06 26.36 6.78 26.39 19.17 21.71
-AVG 11.74 24.13 8.30 25.56 6.46 24.48 18.36 22.12
-STD 0.404 0.392 0.493 0.525 0.466 1.147 0.776 0.990
-k=15 11.42 23.21 8.31 24.68 7.44 21.5 19.12 20.55
-11.52 24.76 8.93 25.77 8.41 25.05 19.79 21.93
-10.14 24.49 7.84 26.46 7.07 23 18.47 24.1
-11.77 23.49 8.11 23.52 7.89 28.41 20.25 24.11
-11.77 23.34 8.52 26.43 6.96 21.25 20.24 23
-AVG 11.32 23.86 8.34 25.37 7.55 23.84 19.57 22.74
-STD 0.680 0.714 0.414 1.262 0.601 2.968 0.770 1.520
-k=20 13.09 22.95 8.02 - - 24.21 18.57 20.79
-11.51 23.58 9.40 - - 21.31 16.66 20.56
-11.76 23.24 8.30 - - 22.27 17.95 23.95
-11.85 23.94 8.20 - - 21.99 19.25 20.61
-11.95 22.80 9.00 - - 25.61 20.19 23.31
-AVG 12.03 23.30 8.58 - - 23.08 18.52 21.84
-STD 0.614 0.465 0.588 - - 1.779 1.333 1.648
+1 in K Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
+k=3 16.93 7.05 19.39 24.89 6.12 20.61 28.82 35.80
+16.88 7.43 21.30 25.07 6.23 22.40 29.24 35.15
+16.55 6.37 20.01 25.09 6.70 20.58 27.82 32.25
+17.34 7.24 22.95 23.82 5.70 20.96 28.97 33.58
+15.25 7.54 19.71 24.59 5.66 21.69 29.18 35.05
+AVG 16.59 7.13 20.67 24.69 6.08 21.25 28.81 34.37
+STD 0.800 0.462 1.466 0.527 0.427 0.784 0.576 1.435
+k=5 17.55 9.72 19.33 24.35 5.63 24.47 27.78 35.70
+15.90 10.50 18.26 24.85 7.20 22.47 29.19 36.12
+15.54 11.02 15.34 24.03 6.12 21.93 30.06 33.99
+16.46 11.55 20.07 23.60 5.90 22.01 28.95 34.39
+16.22 10.36 19.00 23.37 5.86 23.01 28.13 34.48
+AVG 16.33 10.63 18.40 24.04 6.14 22.78 28.82 34.94
+STD 0.763 0.692 1.830 0.591 0.616 1.039 0.901 0.920
+k=7 14.37 12.70 17.83 23.93 7.25 24.52 29.94 34.47
+14.28 12.09 16.69 24.10 6.11 23.09 28.13 34.93
+15.18 10.95 18.25 23.60 6.09 24.05 29.74 32.26
+15.65 11.51 17.68 23.81 6.04 23.47 30.29 35.47
+16.17 11.99 18.22 22.65 5.78 23.77 31.00 35.81
+AVG 15.13 11.85 17.73 23.62 6.25 23.78 29.82 34.59
+STD 0.815 0.657 0.633 0.571 0.572 0.546 1.060 1.398
+k=10 16.04 12.20 15.97 22.10 6.06 21.83 31.15 34.62
+14.90 12.62 15.92 22.78 6.11 22.11 30.38 35.07
+15.24 13.65 16.39 23.42 5.64 24.25 30.47 35.48
+15.55 12.46 16.59 23.37 5.39 22.95 32.16 34.23
+15.56 12.26 17.49 23.85 7.08 25.03 29.40 33.65
+AVG 15.46 12.64 16.47 23.10 6.06 23.23 30.71 34.61
+STD 0.423 0.590 0.635 0.678 0.646 1.376 1.022 0.713
+k=15 15.43 12.39 17.70 26.02 5.51 25.42 31.44 34.84
+15.26 13.18 16.60 25.30 7.23 24.87 31.85 35.15
+15.22 12.76 14.40 25.35 5.49 24.45 30.32 34.60
+15.32 12.57 15.20 25.10 5.72 24.37 31.54 35.83
+14.91 12.79 16.12 26.04 5.76 24.90 32.22 35.69
+AVG 15.23 12.74 16.00 25.56 5.94 24.80 31.47 35.22
+STD 0.195 0.295 1.271 0.437 0.730 0.420 0.713 0.531
+k=20 16.24 12.60 15.07 - - 24.29 31.76 36.16
+15.03 13.46 14.38 - - 23.72 31.80 34.58
+16.00 13.64 15.57 - - 23.43 30.94 36.53
+15.85 13.42 15.13 - - 24.37 31.47 34.87
+15.40 13.03 14.35 - - 24.09 31.84 35.40
+AVG 15.70 13.23 14.90 - - 23.98 31.56 35.51
+STD 0.486 0.417 0.525 - - 0.397 0.377 0.829
 Table 16: 1-in-K ROUGE scores for reference. Please note that ROUGE score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.
+Appendix F Qualitative Error Case Analysis
+Providing the ground-truth clue images effectively makes the problem a standard VQA task, which is considered to be “solved” for the state-of-the-art proprietary models.
+However, our experiment results shown that their improvements over zero-shot baseline is marginal.
+We present a few typical error cases from proprietary models under GT clue setting in Table 17.
+Question: When resting, Acadian Hairstreak (scientific name: Satyrium acadica) usually fold its wings and the abodmen is hidden; what color is its abdomen?
+Ground Truth Answer: [white, light grey]
+Gemini’s Answer: In the provided image, the abdomen of the Acadian Hairstreak is obscured by its folded wings, making its color not visible.
+Question: What color is the stamen tip inside the bell-shaped flower of western moss-heather (scientific name: Cassiope mertensiana)? (Explanation: opening of the flower usually points downward)
+Ground Truth Answer: light yellow
+Gemini’s Answer: Red
+Table 17: Sample cases that Gemini failed to answer even with ground-truth clue image.
+Fine-grained details in the image of area of few pixels remain challenging even for the state-of-the art proprietary models.